### v1

# Robust Adaptive Discrete-Time Control Barrier Certificate

Abstract
This work develops a robust adaptive control strategy for discrete-time systems using Control Barrier Functions (CBFs) to ensure safety under parametric model uncertainty and disturbances. A key contribution of this work is establishing a barrier function certificate in discrete time for general online parameter estimation algorithms. This barrier function certificate guarantees positive invariance of the safe set despite disturbances and parametric uncertainty without access to the true system parameters. In addition, real-time implementation and inherent robustness guarantees are provided. Our approach demonstrates that, using the proposed robust adaptive CBF framework, the parameter estimation module can be designed separately from the CBF-based safety filter, simplifying the development of safe adaptive controllers for discrete-time systems. The resulting safety filter guarantees that the system remains within the safe set while adapting to model uncertainties, making it a promising strategy for real-world applications involving discrete-time safety-critical systems.
keywords:
robust adaptive control; data-driven safety; discrete-time control barrier function; online learning
thanks: This paper is part of a project that has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101018826 - CLariNet), and it was not presented at any IFAC meeting. Corresponding author: Changrui Liu. Tel. +31 15 27 82292.
1 Introduction
Ensuring the safety of dynamical systems is a fundamental requirement in modern control applications, particularly in domains where violations of safety constraints can lead to undesirable outcomes [4], [36]. Safety-critical control focuses on the synthesis of controllers guaranteeing that the system remains within a safe set. Among various approaches (e.g., Hamilton-Jacobi reachability analysis [5], penalty functions [33], and Model Predictive Control (MPC) [20], [30]), Control Barrier Functions (CBFs) have emerged as a powerful and systematic tool to enforce safety for both continuous-time [4] and discrete-time systems [2]. CBF-based control methods allow the formal encoding of safety constraints, and they have found wide-ranging applications, including autonomous driving [8], robotics [16], and power systems [41], demonstrating their versatility and practical relevance in safety-critical domains.
Despite the recent surge in learning-based approaches for safety-critical control [35], model-based methods using CBFs remain essential due to their analytical tractability, real-time implementability, and formal safety guarantees [4], [43]. These approaches allow for rigorous controller design based on dynamical models [2], [4], [46], making them highly attractive for safety-critical applications. However, a fundamental challenge in deploying model-based CBFs is the presence of model uncertainty, which can compromise safety if not properly addressed [43]. To mitigate this, various extensions have been developed that incorporate robust and/or adaptive mechanisms against disturbances [12], [13], [18], [48] or parametric uncertainty [24], [25], [34], [38], [44]. Specifically, for safe control under disturbances, verifiable certificates have been established to ensure safety with high probability for unbounded stochastic disturbances [12], [13] and guaranteed safety for bounded disturbances [18], [48]. On the other hand, recent results on adaptive safety under parametric model mismatch have been obtained using CBFs with online parameter updates, employing either Lyapunov-based methods [24], [25], [38], [44] or batched least-squares methods [34]. For an in-depth overview, we refer the reader to [43].
Although recent developments are promising, to the best of the authors’ knowledge, existing results on adaptive safety for handling parametric uncertainty are limited to continuous-time settings [24], [25], [34], [38], [44], while the discrete-time adaptive case remains unexplored. Given the widespread use of digital controllers and discrete-time formulations in practice—such as in MPC and reinforcement learning [42]—it is critical to develop adaptive safety control within a Discrete-Time CBF (DT-CBF) framework to bridge this gap. Moreover, such discrete-time frameworks are more compatible with real-world applications with discrete-time sensing and actuation [12], [32].
Since time derivatives are not available in discrete time, the main challenge lies in establishing all results using difference equations and inequalities. Moreover, a causal synthesis of the online safe controller must be ensured such that no future data is required for computing the control input., which is an important aspect that is missing when continuous-time models are used [25], [34]. On the other hand, Lyapunov-based adaptation laws often lack guarantees on the boundedness of the parameter estimates [44] or convergence to the true parameters [25], making estimation-based approaches more appealing [10], [44]. Since many simple and effective estimation or identification methods (e.g., least-squares techniques) require access to sampled data [23], a DT-CBF framework is naturally better suited for leveraging these methods. In this paper, we address this theoretical gap by developing a DT-CBF framework for adaptive safety.
Our approach extends model-based safe adaptive control to a broader class of systems—particularly those implemented in embedded or sampled-data settings—and lays the groundwork for online safe control in real-world systems with uncertain models. Specifically, the contributions of this work can be summarized as follows:
  • We develop a robust adaptive Control Barrier Certificate (CBC) for input-affine discrete-time systems subject to both bounded disturbances and parametric model mismatch. To the best of our knowledge, this is the first result on adaptive safety using DT-CBF.
  • The proposed certificate accommodates general online adaptation laws that are decoupled from the CBF-based controller synthesis, enabling a modular control architecture in the spirit of [9]. This new insight—similar to the separation principle in the control and estimation for linear systems [39]—has not been revealed in existing results [25], [34].
  • The inherent robustness111 In this paper, inherent robustness refers to steering the state towards a safe set when the current state is unsafe [46, Section 2.2], which should not be confused with robustness against model uncertainty [15], [25]. of the proposed robust adaptive control approach using DT-CBF is rigorously analyzed, extending existing robustness guarantees for continuous-time systems with time-invariant control barrier certificates [46].
The rest of the paper is organized as follows. Section 2 provides preliminaries. Section 3 introduces the robust adaptive CBC using DT-CBFs. A simple online estimation algorithm based on Recursive Least Squares (RLS) is briefly given in Section 4. Inherent robustness guarantees are discussed in Section 5. A simulation example is presented in Section 6, and Section 7 concludes the paper and outlines future research directions.
2 Preliminaries
This section introduces robust DT-CBFs, which constitute a natural and modest extension of continuous-time CBFs to discrete-time systems with bounded disturbances, building on prior work on continuous-time systems with bounded disturbances [18] and discrete-time systems without disturbances [2].
2.1 Notation
The sets of real and natural numbers are denoted, respectively, by $\mathbb{R}$ and $\mathbb{N}$. The set of extended natural numbers is $\overline{\mathbb{N}}=\mathbb{N}\cup\{+\infty\}$, and $\mathbb{I}_{[a:b]}:=\overline{\mathbb{N}}\cap[a,b]$. The matrix (vector) $p$-norm is denoted by $\|\cdot\|_{p}$ with $p\in\{1,2\}$, and $\|\cdot\|$ represents the $2$-norm by default. The $i$-th element of a vector $x$ is denoted by $x[i]$. The zero and one vectors of size $n$ are denoted by $\mathbf{0}_{n}$ and $\mathbf{1}_{n}$, respectively. A continuous function $\alpha:\mathbb{R}\to\mathbb{R}$ is said to be an extended class $\mathcal{K}_{\infty}$ function (denoted as $\alpha\in\mathcal{K}_{\mathrm{e},\infty}$) if $\alpha(0)=0$, $\alpha$ is strictly monotonically increasing, $\lim_{r\to-\infty}\alpha(r)=$-$\infty$, and $\lim_{r\to+\infty}\alpha(r)=+\infty$ [38, Definition 7]. The class of functions $\beta:\mathbb{R}\to\mathbb{R}$ that are sublinear w.r.t. identity (i.e., $\beta(r)<r$ for all $r>0$) is denoted by $\mathcal{C}_{\mathrm{SL}}$. The projection operator $\Pi_{\mathcal{X}}(\cdot)$ and distance operator $\mathrm{dist}_{\mathcal{X}}(\cdot)$ w.r.t. a set $\mathcal{X}\subseteq\mathbb{R}^{n}$ are defined, respectively, as $\Pi_{\mathcal{X}}(\bar{x})=\arg\min_{x\in\mathcal{X}}\|x-\bar{x}\|$ and $\mathrm{dist}_{\mathcal{X}}(\bar{x})=\min_{x\in\mathcal{X}}\|x-\bar{x}\|$ for any $\bar{x}\in\mathbb{R}^{n}$. The minimum eigenvalue of a symmetric positive definite matrix $\Gamma$ is denoted by $\lambda_{\min}(\Gamma)$.
2.2 System description
Consider a class of input-affine discrete-time systems described by the following model:
$x_{t+1}$ $=f(x_{t},u_{t};\theta)+w_{t}$
$=f_{\mathrm{d}}(x_{t})-[\bm{\phi}(x_{t})]^{\top}\theta+g(x_{t})u_{t}+w_{t},\vskip-5.69046pt$ (1)
where $x_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ and $u_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ are the state and input at time step $t$ with $\mathcal{X}$ being compact and $\mathcal{U}$ being closed, $f_{\mathrm{d}}:\mathcal{X}\to\mathbb{R}^{n}$ describes the nominal drift [^2] Although we consider discrete-time systems, we adopt the term “drift” from the “drift vector field” used in nonlinear control of input-affine continuous-time systems [29]., $g:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ characterizes the input coupling, and $\bm{\phi}(x_{t})]^{\top}\theta$ represents the additional drift term with $\bm{\phi}:\mathbb{R}^{n}\to\mathbb{R}^{q\times n}$ being a continuous known kernel (i.e., regressor) and $\theta\in\Theta$ being an unknown parameter. The considered model class in (2.2) can be used in many real-world applications, e.g., inverted pendulum [27], throttle control [37] and Adaptive Cruise Control (ACC) [38] in automotive systems. In addition, the disturbance $w_{t}$ lies in a known polytope given by
$\mathcal{W}=\{w\in\mathbb{R}^{n}\mid H_{w}w\leq\mathbf{1}_{n_{w}}\},\vskip-5.69046pt$ (2)
where $H_{w}\in\mathbb{R}^{n_{w}\times n}$ and $n_{w}$ is the number of linear inequalities that define $\mathcal{W}$. It is clear that $\mathbf{0}_{n}$ lies in the interior of $\mathcal{W}$, and we further define $\overline{w}:=\max_{w\in\mathcal{W}}\|w\|$. The true parameter $\theta^{\ast}\in\Theta$ is a constant and the prior known parameter set $\Theta$ is a polytope given by
$\Theta=\{\theta\in\mathbb{R}^{p}\mid H_{\theta}\theta\leq h_{\theta}\},\vskip-5.69046pt$ (3)
where $H_{\theta}\in\mathbb{R}^{n_{\theta}\times q}$ and $h_{\theta}\in\mathbb{R}^{n_{\theta}}$. Given an estimate $\hat{\theta}\in\Theta$, its associated estimation error is defined by
$e_{\theta}:=\hat{\theta}-\theta^{\ast}.\vskip-5.69046pt$ (4)
The error $e_{\theta}$ cannot be computed exactly since $\theta^{\ast}$ is unknown. However, for a specific estimate $\hat{\theta}$, an upper bound of the estimation error can be computed by solving the following non-convex optimization problem:
$\varepsilon_{\theta}(p):=\max_{\theta\in\Theta}\|\theta-\hat{\theta}\|_{p},\vskip-5.69046pt$ (5)
and it holds that [^3] Given $x\in\mathbb{R}^{n}$, $\|x\|_{2}\leq\|x\|_{1}$ [17]. $\|e_{\theta}\|\leq\varepsilon_{\theta}(2)\leq\varepsilon_{\theta}(1)$. Following (2.2), the true system under disturbances is given by
$x_{t+1}=f(x_{t},u_{t};\theta^{\ast})+w_{t},\vskip-5.69046pt$ (6)
assuming access to the true parameter $\theta^{\ast}$.
Remark 1.
The controller design for the model (2.2) does not require the smoothness of the kernel $\bm{\phi}(\cdot)$. Therefore, with special non-smooth kernels, the model (2.2) can also describe PWA dynamics [28], [42], and our results can also be used for safety-critical control of PWA systems.
2.3 Robust safety certificate
Definition 1 (Robust safety).
A set $\mathcal{X}_{\mathrm{s}}\subseteq\mathcal{X}$ is positive invariant for (6) if $x_{0}\in\mathcal{X}_{\mathrm{s}}$ implies that $x_{t}\in\mathcal{X}_{\mathrm{s}}$ for all $t\in\mathbb{N}$ and all $w_{t}\in\mathcal{W}$. In this case, the system (6) is robustly safe w.r.t. $\mathcal{X}_{\mathrm{s}}$, and $\mathcal{X}_{\mathrm{s}}$ is called a safe set.
Characterizing safe sets using CBFs is useful for safe controller design [3], [4]. For discrete-time systems, DT-CBFs are powerful tools to achieve this objective [2], [13]. Specifically, given a continuous function $B:\mathcal{X}\to\mathbb{R}$, its induced safe set is defined using the 0-superlevel set as
$\mathcal{S}_{B}:=\{x\in\mathcal{X}\mid B(x)\geq 0\},$ (7)
and its boundary $\partial\mathcal{S}_{B}$ and interior $\mathrm{int}(\mathcal{S}_{B})$ are defined, respectively, by
$\partial\mathcal{S}_{B}$ $:=\{x\in\mathcal{X}\mid B(x)=0\},$ (8a)
$\mathrm{int}(\mathcal{S}_{B})$ $:=\{x\in\mathcal{X}\mid B(x)>0\}.\vskip-5.69046pt$ (8b)
In addition, to certify safety in the discrete-time setting, an additional continuity assumption is required, which is given as follows:
Assumption 1 (Lipschitz continuity).
The barrier function $B:\mathcal{X}\to\mathbb{R}$ is locally Lipschitz continuous on the state space $\mathcal{X}$, i.e.,
$|B(x^{\prime})-B(x^{\prime\prime})|\leq L_{B}\|x^{\prime}-x^{\prime\prime}\|,\quad\forall x^{\prime},x^{\prime\prime}\in\mathcal{X},\vskip-5.69046pt$ (9)
where $L_{B}\geq 0$ is the Lipschitz constant.
Assumption 1 has been used in the literature on discrete-time CBFs [27], [48], and given that the commonly used CBFs are continuously differentiable [3], [4], Lipschitz continuity directly follows given that $\mathcal{X}$ is compact. Besides, the CBFs used in many safe control applications are Lipschitz continuous [1], [25], [38], [48]. The function $B$ is a DT-CBF if it satisfies additional conditions w.r.t. the system (6). The following definition is adopted from [2] and [13] with modifications that suit discrete-time models with bounded disturbances.
Definition 2.
A continuous function $B:\mathcal{X}\to\mathbb{R}$ is a robust DT-CBF for (6) on $\mathcal{S}_{B}$ if there exists $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$ such that for all $x\in\mathcal{X}$, there exists $u\in\mathcal{U}$ satisfying
$B(f(x,u;\theta^{\ast})+w)-B(x)-L_{B}\overline{w}\geq-\alpha\left(B(x)\right),\vskip-5.69046pt$ (10)
and (10) is called the robust CBC against disturbances.
The barrier certificate (10) also defines the set of robust safe control inputs for $\mathcal{S}_{B}$ as
[…]
where $\theta^{\ast}$ emphasizes that characterizing $\mathcal{U}_{\mathrm{rs}}(x;\theta^{\ast})$ requires knowing the true parameter. The following theorem is a standard result on CBF-based safety guarantees, and the proof is omitted for brevity since it follows the same procedure as in [2].
Theorem 1 (Robust safety).
If a continuous function $B:\mathcal{X}\to\mathbb{R}$ is a robust DT-CBF for (6) on $\mathcal{S}_{B}$, then $\mathcal{U}_{\mathrm{rs}}(x;\theta^{\ast})$ is non-empty for all $x\in\mathcal{X}$. In addition, if a state-feedback control policy $\pi:\mathbb{R}^{n}\to\mathbb{R}^{m}$ satisfies $\pi(x)\in\mathcal{U}_{\mathrm{rs}}(x;\theta^{\ast})$, $\forall x\in\mathcal{X}$, then the closed-loop system $x_{t+1}=f(x_{t},\pi(x_{t});\theta^{\ast})+w_{t}$ is robustly safe w.r.t. $\mathcal{S}_{B}$ against disturbances $w_{t}$.
Remark 2.1.
Contrary to the existing results on robust DT-CBFs with (unbounded) stochastic disturbances [12], [13], this paper focuses on bounded disturbances that can be stochastic [28] or adversarial [1]. Accordingly, Definitions 1 and 2 and Theorem 1 are tailored for robust safety against bounded disturbances only, assuming the true parameter $\theta^{\ast}$ is known to the controller.
Since knowing $\theta^{\ast}$ is a stringent requirement in general, (10) cannot be used to synthesize safe controllers when $\theta^{\ast}$ is unknown. In such cases, a straightforward remedy is to additionally robustify against maximized model mismatch and use a constant nominal $\hat{\theta}$ to build the CBF condition. Accordingly, similar to Theorem 1, a corresponding worst-case robust safety guarantee can be established. However, controllers designed based on such worst-case safety condition may not respond to dangerous system behaviors as fast as its counterpart that knows the true model, and more conservative inputs are then needed to guarantee safety. Considering that parametric uncertainty is epistemic in nature [43] and can be reduced or even eliminated through the acquisition of data, it is thus beneficial to develop an adaptive barrier certificate that can be used for safe control synthesis.
3 Robust adaptive safety
In this section, we introduce robust adaptive safety guarantees with general online parameter adaptation laws. More specifically, a robust adaptive CBC is formulated that only relies on an instantaneous parameter estimate and the associated adaptation law. This proposed certificate ensures robust safety against both parametric model mismatch and disturbances with reduced conservatism due to the adaptation mechanism.
3.1 Online parameter estimation
At each time step $t$, consider a parameter estimation module $\mathtt{EST}$ that can provide a polytopic set estimate $\Theta_{t}\subseteq\Theta$ such that $\theta^{\ast}\in\Theta_{t}$ and a point estimate $\hat{\theta}_{t}\in\Theta_{t}$. Following (4) and (5), the instantaneous estimation error and its associated error bound are defined as follows:
$e_{\theta,t}:=\hat{\theta}_{t}-\theta^{\ast},$ (11a)
$\varepsilon_{\theta,t}(p):=\max_{\theta\in\Theta_{t}}\|\theta-\hat{\theta}_{t}\|_{p}.\vskip-5.69046pt$ (11b)
The module $\mathtt{EST}$ performs a recursive point-estimate update specified by the following abstracted update law:
$\hat{\theta}_{t+1}=\mathcal{T}_{\mathtt{ADP},t}(\hat{\theta}_{t}),\vskip-5.69046pt$ (12)
where $\mathcal{T}_{\mathtt{ADP},t}(\cdot)$ is time-varying since the closed-loop input-state data is collected online and the set estimate $\Theta_{t}$ will also be used to determine the point estimate. Finally, the point-estimate increment is defined by
$\delta_{\theta,t}:=\hat{\theta}_{t+1}-\hat{\theta}_{t}=\mathcal{T}_{\mathtt{ADP},t}(\hat{\theta}_{t})-\hat{\theta}_{t},\vskip-5.69046pt$ (13)
which will be used to build the barrier certificate.
3.2 Robust adaptive control barrier certificate
Given a baseline continuous barrier candidate $B:\mathcal{X}\to\mathbb{R}$ that encodes safety constraints and that defines a safe set $\mathcal{S}_{B}$ as (7), we define the corresponding robust adaptive barrier function by
$B_{\mathrm{ra},t}(x):=B(x)-\frac{1}{2}e_{\theta,t}^{\top}\Gamma^{-1}e_{\theta,t},\vskip-5.69046pt$ (14)
where $\Gamma$ is a symmetric positive definite matrix. Accordingly, the robust adaptive safe set $\mathcal{S}_{\mathrm{ra},t}$ is defined by the 0-superlevel set of $B_{\mathrm{ra},t}(\cdot)$ as
$\mathcal{S}_{\mathrm{ra},t}:=\{x\in\mathcal{X}\mid B_{\mathrm{ra},t}(x)\geq 0\},\vskip-5.69046pt$ (15)
and its boundary $\partial\mathcal{S}_{\mathrm{ra},t}$ and interior $\mathrm{int}(\mathcal{S}_{\mathrm{ra},t})$ are defined similar to (8a) and (8b), respectively. Besides, it is noted that (14) implies that $\mathcal{S}_{\mathrm{ra},t}\subseteq\mathcal{S}$, and $\mathcal{S}_{\mathrm{ra},t}=\mathcal{S}$ if and only if $\hat{\theta}_{t}=\theta^{\ast}$. Since $\mathcal{S}_{\mathrm{ra},t}$ is time-varying, a timed version of the robust safety in Definition 1 is given as follows:
Definition 3.1 (Timed robust safety).
Given $T\in\overline{\mathbb{N}}$, a sequence of sets $\{\mathcal{X}_{\mathrm{s},t}\}^{T}_{t=0}$ satisfying $x_{0}\in\mathcal{X}_{\mathrm{s},0}$ and $\forall t\in\mathbb{I}_{[0:T]}$, $\mathcal{X}_{\mathrm{s},t}\subseteq\mathcal{X}$ is sequentially positively invariant for (6) if $\forall t\in\mathbb{I}_{[0:T-1]}$, $x_{t}\in\mathcal{X}_{\mathrm{s},t}$ implies that $x_{t+1}\in\mathcal{X}_{\mathrm{s},t+1}$ for all $w_{t}\in\mathcal{W}$. In this case, the system (6) is said to be robustly safe w.r.t. $\{\mathcal{X}_{\mathrm{s},t}\}^{T}_{t=0}$. In addition, if $\mathcal{X}_{\mathrm{s},t}=\mathcal{X}_{\mathrm{s}}^{\ast}$, $\forall t\in\mathbb{I}_{[0:T]}$, the system (6) is said to be robustly safe w.r.t. $\mathcal{X}_{\mathrm{s}}^{\ast}$ over $\mathbb{I}_{[0:T-1]}$.
The concept of imposing finite-horizon safety has been used in safe control using CBFs [12], [13]. In fact, the timed safety in Definition 3.1 is more practical and general than its infinite-horizon counterpart since safety usually needs only be ensured over the duration of operation [32] and taking $T=+\infty$ recovers infinite-horizon safety.
Remark 3.2.
In this work, we consider a general barrier function $B(\cdot)$, instead of a parameter-dependent barrier function $B(\cdot;\theta)$ as in [25], [38]. In fact, model-free CBFs that are derived solely based on the safety requirements have been used in many real-world safe control applications, e.g., ACC [25], [38], aircraft pitch control [25], and aerial robotics [48]. This framework also avoids repeatedly constructing barrier functions $B(\cdot;\hat{\theta}_{t})$ online based on $\hat{\theta}_{t}$, which can be computationally expensive [35].
Since $\theta^{\ast}$ is unknown, $e_{\theta,t}$ cannot be computed, and thus $\mathcal{S}_{\mathrm{ra},t}$ cannot be explicitly characterized. However, $\mathcal{S}_{\mathrm{ra},t}$ only serves as a fictitious set that will be used to realize safety characterized by $B(\cdot)$, as will be manifested in Theorem 2. Next, we present the robust adaptive CBC, which certifies $B(\cdot)$ as a barrier function for a system in the presence of bounded disturbances and parameteric uncertainty with adaptation. In addition, this CBC can be used in the safety filter framework [4].
Definition 3.3 (Robust adaptive CBC).
Let Assumption 1 hold. Given $\mathtt{EST}$ and $T\in\overline{\mathbb{N}}$, a continuous function $B:\mathcal{X}\to\mathbb{R}$ is a robust adaptive DT-CBF for (6) and $\mathtt{EST}$ on $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$ if there exists $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$ such that, for all $x\in\mathcal{X}$, there exists $u\in\mathcal{U}$ satisfying
$B\big{(}f(x,u;\hat{\theta}_{t})\big{)}-B(x)-L_{B}\overline{w}-E_{\theta,t}(x)\\ \geq-\alpha\left(B(x)-\frac{(\varepsilon_{\theta,t}(p))^{2}}{2\lambda_{\min}(\Gamma)}\right),\forall t\in\mathbb{I}_{[0:T-1]},\vskip-5.69046pt$ (16)
where $E_{\theta,t}(x)\hskip-2.27626pt=\hskip-2.27626pt\left(L_{B}\|\bm{\phi}(x)\|+\frac{\|\delta_{\theta,t}\|}{\lambda_{\min}(\Gamma)}\right)\varepsilon_{\theta,t}(p)+\frac{\|\delta_{\theta,t}\|^{2}}{2\lambda_{\min}(\Gamma)}$ with $L_{B}$ being the Lipschitz constant in (9) and $\delta_{\theta,t}$ characterized by the estimation module $\mathtt{EST}$. In this case, the condition (16) is called the robust adaptive CBC against disturbances and parametric model mismatch.
The certificate (16) also defines a set of robust adaptive safe control inputs for $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$ as
[…]
where the parametric dependency on $\hat{\theta}_{t}$ indicates that characterizing the robust adaptive safe control set only requires the online estimate $\hat{\theta}_{t}$ without knowing the true $\theta^{\ast}$. In essence, $\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)$ is time-varying and it depends on the estimation module $\mathtt{EST}$. Robust adaptive safety is summarized in the following theorem:
Theorem 2 (Robust adaptive safety).
Let Assumption 1 hold and assume $x_{0}\in\mathcal{S}_{\mathrm{ra},0}$. Given $\mathtt{EST}$ and $T\in\overline{\mathbb{N}}$, if $B$ is a robust adaptive DT-CBF for the system (6) and $\mathtt{EST}$ on $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$, then the set $\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)$ is non-empty for all $x\in\mathcal{X}$ and all $t\in\mathbb{I}_{[0,T-1]}$. In addition, if a sequence of time-varying state-feedback control policies $\pi_{t}:\mathcal{X}\to\mathbb{R}^{m},t\in\mathbb{I}_{[0,T-1]}$ satisfies $\pi_{t}(x)\in\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)$ for all $x\in\mathcal{X}$, then the time-varying closed-loop system
$x_{t+1}=f(x_{t},\pi_{t}(x_{t});\theta^{\ast})+w_{t}\vskip-5.69046pt$ (18)
is robustly safe w.r.t. $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$ and thus robustly safe w.r.t. $\mathcal{S}_{B}$ over $\mathbb{I}_{[0:T]}$ (cf. Definition 3.1) against both parametric model mismatch and disturbances.
The proof of Theorem 2 is given in Appendix A. It is worth noting that the certificate (16) requires only $\hat{\theta}_{t}$ and $\varepsilon_{\theta,t}(p)$, without needing knowledge of $\theta^{\ast}$. This advantage makes (16) more readily applicable for CBF-based controller synthesis. Specifically, given a nominal state-feedback policy $\pi_{\mathrm{nom},t}:\mathcal{X}\to\mathbb{R}^{m}$ at time step $t$, a robust adaptive safe policy $\pi_{\mathrm{ras},t}:\mathcal{X}\to\mathcal{U}$ is implicitly given by the following safety filter:
$\pi_{\mathrm{ras},t}(x;p)=\operatorname*{arg\,min}_{\nu\in\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)}\|\nu-\pi_{\mathrm{nom},t}(x)\|.\vskip-5.69046pt$ (19)
Under $\pi_{\mathrm{ras},t}(\cdot)$, the closed-loop system becomes
$x_{t+1}=f(x_{t},\pi_{\mathrm{ras},t}(x_{t};p);\theta^{\ast})+w_{t}.\vskip-5.69046pt$ (20)
Remark 3.4.
The robust adapitve CBF (14) is built using the true parametric error $e_{\theta,t}$ instead of the set-induced error bound as in [25]. In fact, unlike what has been reported in [25], the invariance property of the safe sets defined by the maximized error bound-based robust adaptive CBFs can only be derived in discrete time, but not in continuous time. More details are given in Appendix B.
3.3 Practical implementation
In general, it is not possible to obtain an analytical expression of the safe policy $\pi_{\mathrm{ras},t}(\cdot;p)$. However, in practice, it suffices to only numerically compute a single safe input $u_{\mathrm{ras},t}(p)=\pi_{\mathrm{ras},t}(x_{t};p)$ by solving (19) for a specific $x_{t}$ at time step $t$. Efficient computation of $u_{\mathrm{ras},t}(p)$ requires fast characterization of $\mathcal{U}_{\mathrm{ras}}(x_{t};\hat{\theta}_{t},p)$ by establishing the inequality (16), which depends on $x_{t}$, $\hat{\theta}_{t}$, and $\varepsilon_{\theta,t}(p)$. However, computing $\varepsilon_{\theta,t}(2)$ requires solving a non-convex optimization problem. An alternative is using $\varepsilon_{\theta,t}(1)$, since then the optimization problem in (11b) can be reformulated as a linear program and solved efficiently [40]. However, since $\varepsilon_{\theta,t}(2)\leq\varepsilon_{\theta,t}(1)$, it follows that $\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},1)\subseteq\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},2)$ for all $x\in\mathcal{X}$, meaning the CBC (16) with $\varepsilon_{\theta,t}(1)$ is more conservative.
Remark 3.5.
Employing $\varepsilon_{\theta,t}(1)$ in (16) results in a more conservative safe controller. Thus, $u_{\mathrm{ras},t}(1)$ is likely to be more suboptimal than $u_{\mathrm{ras},t}(2)$, especially in the context of optimal control when the policy $\pi_{\mathrm{nom},t}$ is (implicitly) given by methods using discrete-time control Lyapunov functions (CLFs) [6] or MPC [20], [21], [30].
In general, the optimization problem of the safety filter (19) is non-convex and cannot be solved efficiently. To address this issue, it is beneficial to construct a concave barrier candidate $B(\cdot)$ such that $\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)$ is convex, and thus computing $u_{\mathrm{ras},t}(p)$ requires solving a convex optimization problem, for which many efficient algorithms with convergence and optimality guarantees exist [7]. Note that concavity of $B(\cdot)$ is frequently assumed for DT-CBFs in the literature [2], [14], [47].
4 Online identification & Causal Synthesis
In this section, following the adaptive control approach for linear systems [26], we instantiate an estimator $\mathtt{EST}$ using the least-squares-based recursive update with set-membership projection [19]. In addition, we present a causal integration of the safety filter and the estimator to guarantee that computing $u_{\mathrm{ras},t}(p)$ does not rely on any information from time step $t+1$ onward.
For $t\geq 1$, both $\hat{\theta}_{t}$ and $\Theta_{t}$ are determined by the information flow $\mathcal{F}_{t-1}:=(x_{0},u_{0},x_{1},u_{1},\dots,x_{t-1})$, and a causal synthesis means computing $u_{t}$ based on $\mathcal{F}_{t}$. The initial estimate satisfies $\hat{\theta}_{0}\in\Theta_{0}$ with $\Theta_{0}=\Theta$. Given $x_{t}$, the residual is defined by $r_{t}:=x_{t}-f_{\mathrm{d}}(x_{t-1})-g(x_{t-1})u_{t-1}$, and it holds that $r_{t}+[\bm{\phi}(x_{t-1})]^{\top}\theta^{\ast}\in\mathcal{W}$, which means $\theta^{\ast}\in\Delta_{\theta,t}$ with the non-falsified set $\Delta_{\theta,t}$ given by
$\Delta_{\theta,t}=\left\{\theta\in\mathbb{R}^{p}\mid H_{w}[\bm{\phi}(x_{t-1})]^{\top}\theta\leq\mathbf{1}-H_{w}r_{t}\right\}.\vskip-5.69046pt$ (21)
The set estimate is updated as
$\Theta_{t+1}=\Theta_{t}\cap\Delta_{\theta,t}.\vskip-5.69046pt$ (22)
For the point estimate, a prior is first computed based on the RLS estimation as
$\hat{\theta}_{t+1}^{[\mathrm{p}]}=\hat{\theta}_{t}-\mu_{t}\bm{\phi}(x_{t-1})\left(x_{t}-f(x_{t-1},u_{t-1};\hat{\theta}_{t})\right),\vskip-5.69046pt$ (23)
where $\mu_{t}\in(0,[\|\bm{\phi}(x_{t-1})\|^{2}]^{-1})$ is the learning rate. The final updated estimate $\hat{\theta}_{t+1}$ is computed by set-membership projection as
$\hat{\theta}_{t+1}=\Pi_{\Theta_{t+1}}(\hat{\theta}_{t+1}^{[\mathrm{p}]}),\vskip-5.69046pt$ (24)
and $\delta_{\theta,t}$ in (13) can be obtained accordingly. It is clear that $\delta_{\theta,t}$ is determined by $\mathcal{F}_{t}$, and therefore $u_{t}$, as computed as in (19), also depends on $\mathcal{F}_{t}$. In short, causality is ensured by synchronizing the estimates $\hat{\theta}_{t}$ and $\Theta_{t}$ with the one-step-delayed information flow $\mathcal{F}_{t-1}$.
Algorithm 1 Robust Adaptive Safe Control
$x_{0}$, $\hat{\theta}_{0}$, $\Theta$, $\overline{w}$, $L_{B}$, $\Gamma$, $\{\pi_{\mathrm{nom},t}\}^{T-1}_{t=0}$
$\Theta_{0}\leftarrow\Theta$, $\mu_{0}\leftarrow 0$, $\Delta_{\theta,0}\leftarrow\mathbb{R}^{p}$
for $t=0$ to T1T-1 do
  Compute $\varepsilon_{\theta,t}(p)$ $\triangleright$ (11b)
  Compute $\hat{\theta}_{t+1}$ and $\delta_{\theta,t}$ $\triangleright$ (21)–(24), (13)
  Determine $\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)$ $\triangleright$ (16)
  Compute $u_{t}$ using the safety filter $\triangleright$ (19)
  Obtain $x_{t+1}$ by applying $u_{t}$
end for
Algorithm 1 outlines the core steps of the proposed robust adaptive safe control with online estimation. The required inputs are the initial state $x_{0}$, initial estimates $\hat{\theta}_{0}$ and $\Theta$, constant parameters $\overline{w}$, $L_{B}$, and $\Gamma$, along with nominal policies $\{\pi_{\mathrm{nom},t}\}^{T-1}_{t=0}$. At time step $t$, $\hat{\theta}_{t}$ and $\delta_{\theta,t}$ are first computed to define $\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)$, and $\hat{\theta}_{t+1}$ and $\Theta_{t+1}$ will be used at $t+1$. Given $\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)$, a safe $u_{t}$ is then computed using the safety filter, and the system state evolves to $x_{t+1}$. It is also worth noting that, under Algorithm 1, we have $\Theta_{1}=\Theta_{0}$ and $\hat{\theta}_{1}=\hat{\theta}_{0}$, meaning the non-trivial update starts from $t=1$.
Remark 4.1.
The safe control scheme in Algorithm 1 also applies when the $1$-norm error bound $\varepsilon_{\theta,t}(1)$ is used. Moreover, given a model-based baseline policy $\pi_{\mathrm{nom}}$, the nominal policies $\{\pi_{\mathrm{nom},t}\}^{T-1}_{t=0}$ can also be generated by combining $\mathtt{EST}$, following the certainty-equivalence principle in adaptive control with online estimation [22], [26].
5 Inherent robustness
Inspired by [46], we discuss the inherent robustness property of the CBC in (16). Specifically, when $x_{t_{0}}$ is only safe but not robustly safe (i.e., $x_{t_{0}}\in\mathcal{S}\setminus\mathcal{S}_{\mathrm{ra},t_{0}}$) or even unsafe (i.e., $x_{t_{0}}\notin\mathcal{S}$) for some $t_{0}\in\mathbb{N}$ due to unexpected perturbations [46] or poor parameter estimates, inherent robustness ensures that the safe inputs derived from the safety filter (19) based on the CBC (16) are able to steer the state towards the sequence of robust adaptive safe sets $\{\mathcal{S}_{\mathrm{ra},t}\}^{\infty}_{t=t_{0}}$ asymptotically, i.e., $x_{t}\in\mathcal{S}_{\mathrm{ra},t}$ when $t\to\infty$. This asymptotic convergence w.r.t. a sequence of sets is formally given as follows:
Definition 5.1.
Assume that $\{\mathcal{X}_{t}\}^{\infty}_{t=0}$ with $\mathcal{X}_{t}\subseteq\overline{\mathcal{X}}$ is sequentially positively invariant (cf. Definition 3.1) for a general time-varying system $x_{t+1}=h_{t}(x_{t})+w_{t}$ for all $w_{t}\in\mathcal{W}$, where $h_{t}:\mathcal{X}\to\mathbb{R}^{n}$ is continuous for all $t\in\overline{\mathbb{N}}$. Then $x_{t}$ converges to $\{\mathcal{X}_{t}\}^{\infty}_{t=0}$ asymptotically on $\overline{\mathcal{X}}$ if $\lim_{t\to\infty}\mathrm{dist}_{\mathcal{X}_{t}}(x_{t})=0$ for all $x_{0}\in\overline{\mathcal{X}}$.
Following (15), we define the corresponding time-varying energy functions on $\mathcal{X}$ as follows:
$V_{t}(x)=\begin{cases}0&\text{if }x\in\mathcal{S}_{\mathrm{ra},t}\\ -B_{\mathrm{ra},t}(x)&\text{if }x\in\mathcal{X}\setminus\mathcal{S}_{\mathrm{ra},t}\end{cases},\quad\forall t\in\overline{\mathbb{N}}.\vskip-5.69046pt$ (25)
It is obvious that $\forall t\in\overline{\mathbb{N}},V_{t}(x)>0$ for $x\in\mathcal{X}\setminus\mathcal{S}_{\mathrm{ra},t}$. Given that $\mathcal{S}_{\mathrm{ra},t}$ is sequentially positively invariant (see Appendix A), $V_{t}$ can be used to show that $x_{t}$ converges to $\{\mathcal{S}_{\mathrm{ra},t}\}^{\infty}_{t=0}$ asymptotically if the input $u_{t}$ satisfies (16), which is stated in the following theorem.
Theorem 3 (Inherent robustness guarantee).
Let Assumption 1 hold. If $B:\mathcal{X}\to\mathbb{R}$ is a robust adaptive DT-CBF for (6) and $\mathtt{EST}$ for $T=\infty$ (cf. Definition 3.3), then $x_{t}$ converges to $\{\mathcal{S}_{\mathrm{ra},t}\}^{\infty}_{t=0}$ asymptotically on $\mathcal{X}$ (cf. Definition 5.1) for the system (6) with $u_{t}\in\mathcal{U}_{\mathrm{ras}}(x_{t};\hat{\theta}_{t},p)$.
The proof of Theorem 3 is provided in Appendix C. Theorem 3 directly implies that a sequence of inputs certified by (16) ensure that $x_{t}\in\mathcal{S}_{\mathrm{ra},t}$ when $t\to\infty$. Moreover, although $\mathcal{S}_{\mathrm{ra},t}$ does not generally converge to $\mathcal{S}$, the asymptotic convergence of $\{\mathcal{S}_{\mathrm{ra},t}\}^{\infty}_{t=0}$ is not affected. As discussed in Remark 3.4, maximized error bound-based CBFs [25] can also be used to establish safety, though their induced CBC is more conservative than (16). However, the maximized error bound can exhibit a monotonicity property [25], which can cause the resulting sequence of safe sets to converge to $\mathcal{S}$ under mild conditions. In this case, the inherent robustness property also holds, and more details are given in Appendix C.
6 Application Example
In this section, we consider the ACC problem, which is frequently used as a benchmark in safe control using CBFs [4], [24], [38], [45]. The continuous-time model of the ego vehicle is given by
[…]
where $v$ and $d$ represent the velocity of the ego vehicle and the distance between the ego vehicle and the front vehicle, respectively. The parameter $M$ denotes the effective mass of the ego vehicle. The nominal velocity of the front vehicle, denoted by $v_{\mathrm{f}}$, is unknown but constrained within a known interval $v_{\text{lb}},v_{\text{ub}}]$. The control input $u$ represents the traction force applied by the ego vehicle, while the resistance force $F_{\mathrm{r}}(v)$ accounts for drag effects and is modeled as $F_{\mathrm{r}}(v)=F_{\text{roll}}+\mu_{\text{vis}}v+\mu_{\text{aero}}v^{2}$ for $v>0$, where $F_{\text{roll}}$ is the Coloumb friction force, $\mu_{\text{vis}}$ is the viscous force coefficient, and $\mu_{\text{aero}}$ is the aerodynamic drag coefficient [38], [45]. Both $F_{\text{roll}}$ and $\mu_{\text{vis}}$ are determined empirically offline [45], whereas $\mu_{\text{aero}}$ is identified online. The disturbance $w_{v}$ is included to capture velocity changes due to unmodeled resistance forces (e.g., wind effects), and $w_{d}$ characterizes stochastic variations of $v_{\mathrm{f}}$. These disturbances are bounded as $|w_{v}|\leq\overline{w}_{v}$ and $|w_{d}|\leq\overline{w}_{d}$, and the aerodynamic coefficient satisfies $\mu_{\text{aero}}\in[\mu^{-},\mu^{+}]$, where $\overline{w}_{v}$, $\overline{w}_{d}$, $\mu^{-}$, and $\mu^{+}$ are known constants. For simplicity and real-time computational efficiency [11], [37], forward Euler discretization is applied to obtain the discrete-time model as
[…]
where $\Delta t=0.1[$\mathrm{s}$\$] is the sampling period. The discretized ACC model (27) is consistent with (2.2) with box-constrained $w_{t}=[w_{v,t}\Delta t,w_{d,t}\Delta t]^{\top}$ and $\theta=[\mu_{\text{aero}},v_{\mathrm{f}}]^{\top}$ such that (2) and (3) are satisfied. The safety constraint is $d\geq 1.8v$ [38], and the barrier function $B$ is chosen as B(v,d)=d1.8vaB(v,d)=d-1.8v-a, where $a\geq 0$ is user-specified safety margin. The nominal controller is an adaptive MPC controller equiped with $\mathtt{EST}$ using the certainty-equivalence principle [20], [22], [26], and the objective of the MPC is to track a reference velocity $v_{\text{ref}}=30[$\mathrm{m}\mathrm{/}\mathrm{s}$\$].
Figure 1: $\mathtt{EST}$ estimation performance. The shaded area is the min-max envelope encompassing different disturbance realizations.
The simulations [^4] The full codebase and specific implementations are available at: https://github.com/lcrekko/ra_cbf_mpc. are conducted using Python 3.12. The performance of the online parameter estimation module $\mathtt{EST}$ is illustrated in Figure 1, which demonstrates that the true parameters are accurately recovered over time. The results related to safety are presented in Figure 2. As shown in the middle subplot, the proposed robust adaptive CBF-based safety filter ensures constraint satisfaction even in the presence of disturbances and model mismatch. This is achieved by actively modifying unsafe control inputs when safety is at risk, as evidenced in the bottom subplot of the same figure.
Based on Figure 2, the need for an explicit safety mechanism becomes apparent when comparing with the baseline adaptive MPC controller, which fails to enforce the safety constraint. Although the robust CBF controller also maintains safety, it reacts more slowly to critical situations compared to the robust adaptive CBF approach. As a result, from approximately $7$ seconds onward, stronger braking actions–i.e., more conservative control inputs–are required to uphold safety.
Figure 2: Comparison of adaptive MPC with robust adaptive CBF-based safety filter (aMPC-raCBF), MPC with worst-case robust CBF-based safety filter (MPC-rCBF), and adaptive MPC (aMPC). The shaded area is the min-max envelope encompassing different disturbance realizations. Zoomed-in subplot is provided for each subfigure.
7 Conclusions
In this paper, we have developed a safe control method for discrete-time systems using robust adaptive discrete-time control barrier functions (CBFs), considering both parametric model uncertainty and disturbances. Unlike traditional approaches in continuous-time adaptive safe control design that tightly couple adaptation and control, our framework supports a modular architecture where the parameter learning algorithm operates independently of the safety filter. This separation simplifies safe control design and supports greater flexibility in implementation. In addition, the proposed method provides rigorous guarantees of safety without requiring exact model knowledge, which is particularly valuable for safety-critical applications where online adaptation is necessary. Future research directions include suboptimality analysis of safe control using CBFs and extension to distributed safe control of multi-agent systems.
Appendix A Proof of Theorem 2
{pf}
It suffices to prove robust safety through positive invariance. Considering $B_{\mathrm{ra},t+1}(x_{t+1})-B_{\mathrm{ra},t}(x_{t})$, which characterizes the incremental value of the barrier function, it holds that
$B_{\mathrm{ra},t+1}(x_{t+1})-B_{\mathrm{ra},t}(x_{t})$
$=$ $B(x_{t+1})-B(x_{t})-\frac{1}{2}(e_{\theta,t+1}\hskip-2.27626pt-\hskip-2.27626pte_{\theta,t})^{\top}\Gamma^{-1}(e_{\theta,t+1}\hskip-2.27626pt+\hskip-2.27626pte_{\theta,t})$
[…]
$\geq$ $B(x_{t+1})-B(x_{t})-\frac{\|\delta_{\theta,t}\|^{2}+2\|\delta_{\theta,t}\|\varepsilon_{\theta,t}(p)}{2\lambda_{\min}(\Gamma)},\vskip-14.22636pt$ (28)
where $B(x_{t+1})-B(x_{t})$ can be lower bounded as
$B(x_{t+1})-B(x_{t})$
$=$ $B(f(x_{t},\pi_{t}(x_{t});\theta^{\ast})+w_{t})-B(x_{t})$
[…]
By substituting the above equation into (A) and using the fact that $\pi_{t}(x_{t})$ satisfies (16), we have
$B_{\mathrm{ra},t+1}(x_{t+1})-B_{\mathrm{ra},t}(x_{t})\\ \geq-\alpha\left(B(x_{t})-\frac{(\varepsilon_{\theta,t}(p))^{2}}{2\lambda_{\min}(\Gamma)}\right)\geq-\alpha(B_{\mathrm{ra},t}(x_{t})).\vskip-5.69046pt$ (29)
Based on (29), if $x_{t}\in\mathcal{S}_{\mathrm{ra},t}$, i.e., $B_{\mathrm{ra},t}(x_{t})\geq 0$, then $B_{\mathrm{ra},t+1}(x_{t+1})\geq B_{\mathrm{ra},t}(x_{t})-\alpha(B_{\mathrm{ra},t}(x_{t}))\geq 0$ since $\alpha\in\mathcal{C}_{\mathrm{SL}}$, which implies $x_{t+1}\in\mathcal{S}_{\mathrm{ra},t+1}$. Thus, $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$ is sequentially positive invariant for the closed-loop system (18) given $\pi_{t}(x)\in\mathcal{U}_{\mathrm{ras}}(x;\hat{\theta}_{t},p)$. Moreover, noting that $B_{\mathrm{ra},t}(x)\geq 0\Longrightarrow B(x)\geq 0$, (29) indicates that $\mathcal{S}_{B}$ is positive invariant for (18) over $\mathbb{I}_{[0:T]}$.       ∎
Appendix B Robustified CBFs with error bounds
The true error $e_{\theta,t}$ is used to construct $B_{\mathrm{ra},t}$ in (14) to achieve adaptive safety, which is similar to [38]. Although safety can be established, $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$ cannot be characterized since $e_{\theta,t}$ is unknown. Considering that the ultimate goal is to achieve safety w.r.t. $\mathcal{S}_{B}$, which encodes safety [25], [38], the proposed method using $B_{\mathrm{ra},t}$ defined in (14) is still valid since $\mathcal{S}_{\mathrm{ra},t}\subseteq\mathcal{S}_{B}$.
An alternative is using the set-induced error bound $\eta_{\theta,t}$ satisfying $|e_{\theta,t}|\leq\eta_{\theta,t}$ to construct robustified CBFs [25]. Specifically, for all $i\in\mathbb{I}_{[1:q]}$, $\eta_{\theta,t}[i]$ can be computed as
$\eta_{\theta,t}[i]=\max_{\theta\in\Theta_{t}}\theta[i]-\min_{\theta\in\Theta_{t}}\theta[i].\vskip-5.69046pt$ (30)
The bound in (30) is conservative since it is solely computed using $\Theta_{t}$. Besides, the $\max$ ($\min$) operation prevents obtaining an analytical expression of its time derivative, and proving set invariance in continuous-time settings has to resort to the true error instead of the bound itself that is used to construct the CBF (see the proof of [25, Theorem 2]). The consequence is that, in the continuous-time setting, it is impossible [^5] Noting that, in the proof of [25, Theorem 2], $h\geq h_{r}-\frac{1}{2}\tilde{\vartheta}^{\top}\Gamma^{-1}\tilde{\vartheta}$ and $h\geq 0$ cannot lead to $h_{r}-\frac{1}{2}\tilde{\vartheta}^{\top}\Gamma^{-1}\tilde{\vartheta}\geq 0$. to establish the invariance property of the safe sets defined by the robustified CBFs constructed using the error bound. Luckily, in the discret-time setting, such invaraince property can be established since no derivative is required. Define the error-bound robust adaptive CBF by
$\overline{B}_{\mathrm{ra},t}(x):=B(x)-\frac{1}{2}\eta_{\theta,t}^{\top}\Gamma^{-1}\eta_{\theta,t},\vskip-5.69046pt$ (31)
where $\Gamma$ is a symmetric positive definite matrix. Similar to (15), the safe set corresponding to $\overline{B}_{\mathrm{ra},t}$ is given by
$\overline{\mathcal{S}}_{\mathrm{ra},t}=\{x\in\mathcal{X}\mid\overline{B}_{\mathrm{ra},t}(x)\geq 0\}.\vskip-5.69046pt$ (32)
Following Definition 3.3, a modified version based on (31) and (32) is given as follows:
Definition B.1.
Let Assumption 1 hold. Given $\mathtt{EST}$ and $T\in\overline{\mathbb{N}}$, A continuous function $B:\mathcal{X}\to\mathbb{R}$ is an error-bound robust adaptive DT-CBF for (6) and $\mathtt{EST}$ on $\{\overline{\mathcal{S}}_{\mathrm{ra},t}\}^{T}_{t=0}$ if there exists $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$ such that, for all $x\in\mathcal{X}$, there exists $u\in\mathcal{U}$ satisfying
$B\big{(}f(x,u;\hat{\theta}_{t})\big{)}-B(x)-L_{B}\overline{w}-E_{\eta,t}(x)\\ \geq-\alpha\left(B(x)-\frac{\|\eta_{\theta,t}\|^{2}}{2\lambda_{\min}(\Gamma)}\right),\forall t\in\mathbb{I}_{[0:T-1]}\vskip-5.69046pt$ (33)
where $E_{\eta,t}(x)\hskip-2.27626pt=\hskip-2.27626pt\left(L_{B}\|\bm{\phi}(x)\|+\frac{\|\delta_{\eta,t}\|}{\lambda_{\min}(\Gamma)}\right)\eta_{\theta,t}+\frac{\|\delta_{\eta,t}\|^{2}}{2\lambda_{\min}(\Gamma)}$ with $L_{B}$ being the Lipschitz constant in (9) and $\delta_{\eta,t}=\eta_{\theta,t+1}-\eta_{\theta,t}$. In this case, the condition (33) is called the error-bound robust adaptive CBC against disturbances and parametric model mismatch.
The certificate (33) also defines a set of safe inputs for $\{\overline{\mathcal{S}}_{\mathrm{ra},t}\}^{T}_{t=0}$ as
[…]
Following Theorem 2, a modified safety guarantee that uses the certificate (33) is given as follows:
Theorem 4.
Let Assumption 1 hold and assume $x_{0}\in\overline{\mathcal{S}}_{\mathrm{ra},0}$. Given $\mathtt{EST}$ and $T\in\overline{\mathbb{N}}$, If $B:\mathcal{X}\to\mathbb{R}$ is an error-bound robust adaptive DT-CBF for the system (6) and $\mathtt{EST}$ on $\{\overline{B}_{\mathrm{ra},t}\}^{T}_{t=0}$, then the set $\overline{\mathcal{U}}_{\mathrm{ras}}(x;\hat{\theta}_{t})$ is non-empty for all $x\in\mathcal{X}$ and all $t\in\mathbb{I}_{[0,T-1]}$. In addition, if a sequence of time-varying state-feedback control policies $\overline{\pi}_{t}:\mathcal{X}\to\mathbb{R}^{m},t\in\mathbb{I}_{[0,T-1]}$ satisfies $\overline{\pi}_{t}(x)\in\overline{\mathcal{U}}_{\mathrm{ras}}(x;\hat{\theta}_{t})$ for all $x\in\mathcal{X}$, then the time-varying closed-loop system
$x_{t+1}=f(x_{t},\overline{\pi}_{t}(x_{t});\theta^{\ast})+w_{t}\vskip-5.69046pt$ (35)
is robustly safe w.r.t. $\{\overline{\mathcal{S}}_{\mathrm{ra},t}\}^{T}_{t=0}$ and thus robustly safe w.r.t. $\mathcal{S}_{B}$ over $\mathbb{I}_{[0:T]}$ (cf. Definition 3.1) against both parametric model mismatch and disturbances.
The proof of Theorem 4 is similar to that of Theorem 2 in Appendix A and is therefore omitted for brevity.
Remark B.2.
The advantage of using $\xi_{\theta,t}$ to design the safety certificate is that the safe set $\overline{B}_{\mathrm{ra},t}$ can be explicitly determined. Moreover, it holds that $\eta_{\theta,t+1}\leq\eta_{\theta,t}$ if $\Theta_{t+1}\subseteq\Theta_{t}$, leading to $\overline{\mathcal{S}}_{\mathrm{ra},t}\subseteq\overline{\mathcal{S}}_{\mathrm{ra},t+1}$ (i.e., the safe set is gradually expanding).
Appendix C More on Inherent Robustness
We first prove Theorem 3 using $V_{t}$ defined in (25). The proof is inspired by the proof of [30, Theorem B.13]. {pf} If there exists a $t_{0}\in\mathbb{N}$ such that $x_{t_{0}}\in\mathcal{S}_{\mathrm{ra},t_{0}}$, then $x_{t}\in\mathcal{S}_{\mathrm{ra},t}$ holds for all $t\in\mathbb{I}_{[t_{0}:\infty]}$ by Theorem 2 since $\{B_{\mathrm{ra},t}\}^{\infty}_{t=0}$ is sequentially positive invariant for the system (6) with $u_{t}\in\mathcal{U}_{\mathrm{ras}}(x_{t};\hat{\theta}_{t},p)$. Thus, in this case, $\lim_{t\to\infty}\mathds{1}(x_{t}\in\mathcal{S}_{\mathrm{ra},t})=1$. On the other hand, assume $x_{t}\in\mathcal{X}\setminus\mathcal{S}_{\mathrm{ra},t}$; then we have
[…]
where the last inequality is due to $B_{\mathrm{ra},t}(x_{t})<0$ for $x_{t}\in\mathcal{X}\setminus\mathcal{S}_{\mathrm{ra},t}$ and $\alpha\in\mathcal{K}_{\mathrm{e},\infty}$. Consider the sequence $\{V_{t}(x_{t})\}^{\infty}_{t=0}$; given that $V_{t}(x_{t})\geq 0$ and (36), by the monotone convergence theorem [31], there exists a $V^{\ast}\geq 0$ such that $\lim_{t\to\infty}V_{t}(x_{t})=V^{\ast}$. Equivalently, $\forall\epsilon>0$, $\exists T^{\ast}(\epsilon)$ such that $\forall t\geq T^{\ast}(\epsilon)$, $V_{t}(x_{t})-V^{\ast}<\epsilon$. Next, we prove that $V^{\ast}=0$ by contradiction. Assume $V^{\ast}>0$, and let $\epsilon=-\alpha(-V^{\ast})$. Then, for any $t\geq T^{\ast}(-\alpha(-V^{\ast}))$, it holds that (i) $V^{\ast}\leq V_{t}(x_{t})<V^{\ast}-\alpha(-V^{\ast})$ and (ii) $V^{\ast}\leq V_{t+1}(x_{t+1})<V^{\ast}-\alpha(-V^{\ast})$, implying $V_{t+1}(x_{t+1})-V_{t}(x_{t})>\alpha(-V^{\ast})$. On the other hand, by (36), we have $V_{t+1}(x_{t+1})-V_{t}(x_{t})\leq\alpha(-V^{\ast})$, which leads to a contradiction. Therefore, $\lim_{t\to\infty}V_{t}(x_{t})=V^{\ast}=0$, i.e., $\lim_{t\to\infty}\mathrm{dist}_{\mathcal{S}_{\mathrm{ra},t}}(x_{t})=0$.             ∎
In fact, following the discussion in Appendix B, by constructing a sequence of energy functions using $\overline{\mathcal{S}}_{\mathrm{ra},t}$, it can been shown that the safe control input characterized by the certificate (33) also has the inherent robustness property, which is summarized in the following theorem:
Theorem 5.
Let Assumption 1 hold. If $B:\mathcal{X}\to\mathbb{R}$ is an error-bound robust adaptive DT-CBF for (6) and $\mathtt{EST}$ for $T=+\infty$ (cf. Definition B.1, then $x_{t}$ converges to $\{\overline{\mathcal{S}}_{\mathrm{ra},t}\}^{\infty}_{t=0}$ defined by the robust adaptive CBFs $\{\overline{B}_{\mathrm{ra},t}\}^{\infty}_{t=0}$ asymptotically (cf. Definition 5.1) on $\mathcal{X}$ for the system (6) with $u_{t}\in\overline{\mathcal{U}}_{\mathrm{ras}}(x_{t};\hat{\theta}_{t})$.
Theorem 5 implies that $x_{t}$ will be regulated towards the safe sets $\{\overline{\mathcal{S}}_{\mathrm{ra},t}\}^{\infty}_{t=0}$ asymptotically on $\mathcal{X}$ for the closed-loop system (35). The proof of Theorem 5 closely follows that of Theorem 3 given above and is therefore omitted for brevity.