### v2

# Robust Adaptive Discrete-Time Control Barrier Certificate

Abstract
This work develops a robust adaptive control strategy for discrete-time systems using Control Barrier Functions (CBFs) to ensure safety under parametric model uncertainty and disturbances. A key contribution of this work is establishing a barrier function certificate in discrete time for general online parameter estimation algorithms. This barrier function certificate guarantees positive invariance of the safe set despite disturbances and parametric uncertainty without access to the true system parameters. In addition, real-time implementation and inherent robustness guarantees are provided. The proposed robust adaptive safe control framework demonstrates that the parameter estimation module can be designed separately from the CBF-based safety filter, simplifying the development of safe adaptive controllers for discrete-time systems. The resulting safe control approach guarantees that the system remains within the safe set while adapting to model uncertainties, making it a promising strategy for discrete-time safety-critical systems.
keywords:
robust adaptive control; data-driven safety; discrete-time control barrier function; online learning
thanks: This paper is part of a project that has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 101018826 - CLariNet), and it was not presented at any IFAC meeting. Corresponding author: Changrui Liu. Tel. +31 15 27 82292.
1 Introduction
Safety-critical control aims to guarantee that the system state remains within a safe set, i.e., constraint satisfaction, which is crucial in domains where unsafe system behaviors can lead to undesirable outcomes [3]. Among various approaches (e.g., Hamilton-Jacobi reachability analysis [35] and Model Predictive Control (MPC) [21], [32]), Control Barrier Functions (CBFs) becomes a powerful and systematic tool for both continuous-time [3] and discrete-time systems [2]. CBFs allow formal encoding of safety constraints as control invariant sets [4], and they can be integrated on top of traditional control methods (e.g., control Lyapunov functions [3], [19]) and learning-based approaches (e.g., reinforcement learning [9]) that may not inherently offer safety guarantees. CBFs have many applications, including control of bipedal robots [2], Adaptive Cruise Control (ACC) [3], and vehicle engine control [38], demonstrating its versatility and practical relevance in safety-critical domains.
Despite the recent surge in learning-based approaches in CBFs [37], model-based methods remain essential due to their analytical tractability and formal safety guarantees [3], [40]. Typically, these approaches support modular design through the explicit use of dynamical models [2], [3], [43]. A standard approach is to encode safety constraints through a CBF-based inequality condition on the input, and then to incorporate this condition within a safety filter [12] that minimizes the adjustment of a given nominal input that may otherwise lead to unsafe behaviors (i.e., constraint violations). However, a fundamental challenge in model-based methods is model uncertainty, which can compromise safety [40]. Accordingly, various extensions have developed robust and adaptive mechanisms tailored for CBFs to handle disturbances [10], [11], [15], [45] or parametric uncertainty [25], [36], [39], [41]. For disturbances, high-probability safety [11] and guaranteed safety [15] have been established for unbounded and bounded disturbances, respectively. There are also results on adaptive safety focusing on parametric uncertainty, employing either Lyapunov-based methods [25], [39], or batched least-squares methods [36]. For an in-depth overview, we refer the reader to [40].
Although recent progress has been made on robust and adaptive CBFs in continuous time [25], [36], [39], [41], the discrete-time case remains largely unexplored despite its relevance to MPC and reinforcement learning [9]. Since practical systems rely on sampled sensing and actuation, there is a need for adaptive safe control directly based on Discrete-Time CBFs (DT-CBFs) [10], [34]. The main difficulty is that derivative-based arguments are unavailable for DT-CBFs. In addition, a causal controller must compute inputs without using future information, a feature not discussed in continuous-time methods [25], [36]. Besides, Lyapunov adaptation laws lack general guarantees on bounded updates or convergence [41], [25], motivating estimation-based approaches that better align with sampled-data settings [9]. Since standard estimation tools (e.g., least-squares methods) operate on discrete measurements [24], a DT-CBF framework provides a more suitable foundation for adaptive safety.
Our approach extends model-based safe adaptive control to a broader class of systems, particularly those implemented in embedded or sampled-data settings. The contributions of this work are summarized as follows:
  • Robust adaptive safety guarantees using DT-CBFs for input-affine discrete-time systems under bounded disturbances and parametric mismatch are established, and an online input condition is derived to facilitate their deployment in safety filters.
  • The proposed approach accommodates general parameter estimators that are decoupled from the CBF-based controller synthesis, enabling a modular control architecture in the spirit of [8]. This new insight has not been revealed in existing results [25], [36], [39].
  • The inherent robustness [^1] In this paper, inherent robustness refers to steering the state towards safe sets when the current state is unsafe [43, Section 2.2], which should not be confused with robustness against model uncertainty [12], [25].of the proposed safe control approach is rigorously analyzed, extending existing robustness guarantees for continuous-time systems without model uncertainty [43].
It is further noted that the primary contribution of this paper lies in establishing a general robust adaptive safe control framework using DT-CBFs, rather than in designing CBFs, in developing novel parameter estimation algorithms, or in verifying the existence of barrier functions [31]. The rest of the paper is organized as follows. Section 2 provides preliminaries, and Section 3 introduces robust adaptive DT-CBFs. A simple estimation algorithm is given in Section 4. Inherent robustness guarantees are discussed in Section 5. A simulation example is presented in Section 6, and Section 7 concludes the paper and outlines future research directions.
2 Preliminaries
This section lays the groundwork for the main results. After introducing notations and the system model, we present robust DT-CBFs and clarify how CBFs define safe sets and translate state safety requirements into an input feasibility condition via positive set invariance. In essence, CBFs restrict admissible control inputs to ensure that the state remains within the safe set or is driven back to it if it attempts to leave.
2.1 Notation
The sets of real and natural numbers are denoted, respectively, by $\mathbb{R}$ and $\mathbb{N}$, $\overline{\mathbb{N}}:=\mathbb{N}\cup\{+\infty\}$, and $\mathbb{I}_{[a:b]}:=\overline{\mathbb{N}}\cap[a,b]$. The matrix (vector) $p$-norm is denoted by $\|\cdot\|_{p}$, and $\|\cdot\|$ represents the $2$-norm by default. The $i$-th element of a vector $x$ is denoted by $x[i]$. The zero vector, one vector, and identity matrix of size $n$ are denoted by $\mathbf{0}_{n}$, $\mathbf{1}_{n}$, and $\mathbf{I}_{n}$, respectively. A continuous function $\alpha:\mathbb{R}\to\mathbb{R}$ is said to be an extended class $\mathcal{K}_{\infty}$ function (denoted as $\alpha\in\mathcal{K}_{\mathrm{e},\infty}$) if $\alpha(0)=0$, $\alpha$ is strictly monotonically increasing, $\lim_{r\to-\infty}\alpha(r)=$-$\infty$, and $\lim_{r\to+\infty}\alpha(r)=+\infty$ [39, Definition 7]. The class of functions $\beta:\mathbb{R}\to\mathbb{R}$ that are sublinear w.r.t. identity (i.e., $\beta(r)<r$ for all $r>0$) is denoted by $\mathcal{C}_{\mathrm{SL}}$. The projection operator $\Pi_{\mathcal{X}}(\cdot)$ and distance operator $\mathrm{dist}_{\mathcal{X}}(\cdot)$ w.r.t. a closed set $\mathcal{X}\subseteq\mathbb{R}^{n}$ are defined, respectively, as $\Pi_{\mathcal{X}}(\bar{x})=\arg\min_{x\in\mathcal{X}}\|x-\bar{x}\|$ and $\mathrm{dist}_{\mathcal{X}}(\bar{x})=\min_{x\in\mathcal{X}}\|x-\bar{x}\|$ for any $\bar{x}\in\mathbb{R}^{n}$. The minimum eigenvalue of a symmetric positive definite matrix $\Gamma$ is denoted by $\lambda_{\min}(\Gamma)$.
2.2 System description
Consider a class of input-affine discrete-time systems described by the following model:
$x_{t+1}$ $=f(x_{t},u_{t};\theta)+w_{t}$
$=f_{\mathrm{d}}(x_{t})-[\bm{\phi}(x_{t})]^{\top}\theta+g(x_{t})u_{t}+w_{t},\vskip-8.5359pt$ (1)
where $x_{t}\in\mathcal{X}\subset\mathbb{R}^{n}$ and $u_{t}\in\mathcal{U}\subseteq\mathbb{R}^{m}$ are the state and input at time step $t$ with $\mathcal{X}$ being compact and $\mathcal{U}$ being closed, $f_{\mathrm{d}}:\mathcal{X}\to\mathbb{R}^{n}$ describes the nominal drift [^2] Although we consider discrete-time systems, we adopt the term “drift” from the “drift vector field” used in nonlinear control of input-affine continuous-time systems [30].corresponding to known dynamics (e.g., inertia in robotics [11]) and/or $x_{t}$ itself arising from discretization [11], [29], $g:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ characterizes the input coupling, and $\bm{\phi}(x_{t})]^{\top}\theta$ represents the uncertain drift term with $\bm{\phi}:\mathbb{R}^{n}\to\mathbb{R}^{q\times n}$ being a continuous known kernel (i.e., regressor) and $\theta\in\Theta$ being an unknown parameter. The uncertain drift can be used to model unknown damping [28] or state-coupled disturbances [39]. A compact state space aligns with bounded safe sets [3], and a closed input space guarantees the existence of feasible solutions to the safety filter optimization problem. These assumptions are practically valid for most engineering systems as they are designed to (or due to physical constraints) lie in a bounded operational regime [2], [3], [25]. The state $x_{t}$ is considered to be fully measured without error, and thus the values $f_{\mathrm{d}}(x_{t})$, $\bm{\phi}(x_{t})$, and $g(x_{t})$ can be obtained accurately. In addition, the disturbance $w_{t}$ lies in a known polytope [^3] A polytope is a bounded polyhedron (i.e., an intersection of a finite set of closed half-spaces) [6, Chapter 2.1].given by
$\mathcal{W}=\{w\in\mathbb{R}^{n}\mid H_{w}w\leq\mathbf{1}_{n_{w}}\},\vskip-5.69046pt$ (2)
where $H_{w}\in\mathbb{R}^{n_{w}\times n}$ and $n_{w}\geq n$ is the number of linear inequalities that define $\mathcal{W}$. It is clear that $\mathbf{0}_{n}$ lies in the interior of $\mathcal{W}$, and we further define $\overline{w}:=\max_{w\in\mathcal{W}}\|w\|$. The true parameter $\theta^{\ast}\in\Theta$ is a constant and the prior known parameter set $\Theta$ is a polytope given by
$\Theta=\{\theta\in\mathbb{R}^{q}\mid H_{\theta}\theta\leq h_{\theta}\},\vskip-5.69046pt$ (3)
where $H_{\theta}\in\mathbb{R}^{n_{\theta}\times q}$ with $n_{\theta}\geq q$, and $h_{\theta}\in\mathbb{R}^{n_{\theta}}$. Both the disturbance and the model parameter are considered to lie in polytopes, which is motivated by practical considerations, as many disturbances and parameters are naturally bounded and can be modeled by box-constrained sets [1], [3], [25], [45]. Theoretically, polytopes can also inner-approximate arbitrary compact sets, making this a flexible and effective modeling choice. Given an estimate $\hat{\theta}\in\Theta$, its associated estimation error is defined as
$e_{\theta}:=\hat{\theta}-\theta^{\ast}.\vskip-8.5359pt$ (4)
The estimation error $e_{\theta}$ cannot be computed exactly since $\theta^{\ast}$ is unknown. However, for a specific $\hat{\theta}$, an upper bound of its associated estimation error can be computed by solving the following optimization problem:
$\varepsilon_{\theta}(p):=\max_{\theta\in\Theta}\|\theta-\hat{\theta}\|_{p},\vskip-5.69046pt$ (5)
[…]
$x_{t+1}=f(x_{t},u_{t};\theta^{\ast})+w_{t}.\vskip-5.69046pt$ (6)
The system state evolves according to (6); however, the controller does not have access to the true parameter $\theta^{\ast}$.
Remark 1.
The controller design for the model (2.2) does not require the smoothness of the kernel $\bm{\phi}(\cdot)$; with non-smooth kernels, (2.2) can also represent PWA dynamics [29]. The input-affine structure is standard in nonlinear control [30], and linear dependence on the unknown parameters is powerful in kernel modeling [18]. This modeling choice does not trivialize the safe control problem, and it is widely used in practice (e.g., inverted pendulum control [28], throttle control [38], and adaptive cruise control [39]) and enables direct comparison with existing adaptive safe control methods [25], [39].
2.3 Robust safe control
This subsection introduces robust DT-CBFs, which constitute a natural and modest extension of continuous-time CBFs to discrete-time systems with bounded disturbances, building on prior work on continuous-time systems with bounded disturbances [15] and discrete-time systems without disturbances [2].
Definition 1 (Robust safety).
A set $\mathcal{X}_{\mathrm{s}}\subseteq\mathcal{X}$ is positive invariant for (6) if $x_{0}\in\mathcal{X}_{\mathrm{s}}$ implies that $\forall\{w_{t}\}^{\infty}_{t=0}\in\mathcal{W}^{\infty}$, $x_{t}\in\mathcal{X}_{\mathrm{s}}$ for all $t\in\mathbb{N}$. In this case, the system (6) is robustly safe w.r.t. $\mathcal{X}_{\mathrm{s}}$, and $\mathcal{X}_{\mathrm{s}}$ is called a robust safe set.
Definition 1 considers robustness only against the disturbances, which is valid for the true parameter $\theta^{\ast}$ but may not hold for all parameters $\theta\in\Theta$. Characterizing safe sets using CBFs is useful for safe controller design [3]. For discrete-time systems, DT-CBFs are powerful tools to achieve this objective [2], [11]. Specifically, for a continuous barrier function $B:\mathcal{X}\times\Theta\to\mathbb{R}$, its induced safe set is defined using the 0-superlevel set of $B(\cdot,\theta):\mathcal{X}\to\mathbb{R}$ for all $\theta\in\Theta$ as
$\mathcal{S}_{B,\theta}:=\{x\in\mathcal{X}\mid{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B(x,\theta)}\geq 0\},\vskip-5.69046pt$ (7)
and its boundary $\partial\mathcal{S}_{B,\theta}$ and interior $\mathrm{int}(\mathcal{S}_{B,\theta})$ are defined, respectively, by $\partial\mathcal{S}_{B,\theta}:=\{x\in\mathcal{X}\mid{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B(x,\theta)}=0\}$ and $\mathrm{int}(\mathcal{S}_{B,\theta}):=\{x\in\mathcal{X}\mid{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B(x,\theta)}>0\}$.
Remark 2.
This work adopts barrier functions that depend on model parameter $\theta$ as in [25] and [39] since most barrier functions are model-based [3]. Nonetheless, CBFs that are derived mainly based on the safety requirements without using $\theta$ are also commonly used in many safe control applications, e.g., adaptive cruise control [39], aircraft pitch control [25], and aerial robotics [45].
In addition, a continuity assumption is useful to quantify the perturbation of $B$, which is given as follows:
Assumption 1 (Lipschitz continuity).
The function $B:\mathcal{X}\times\Theta\to\mathbb{R}$ is locally Lipschitz continuous on the joint space $\mathcal{X}\times\Theta$, i.e., $\forall x^{\prime},x^{\prime\prime}\in\mathcal{X},\;\theta^{\prime},\theta^{\prime\prime}\in\Theta$,
[…]
with Lipschitz constants $L_{B,x},L_{B,\theta}\geq 0$.
Lipschitz continuity w.r.t. the state $x$ has been used in the literature on discrete-time CBFs [28], [45]. Moreover, when $f(x,u;\cdot):\Theta\to\mathcal{X}$ is continuously differentiable (see the model in (2.2)), the commonly constructed CBFs $B:\mathcal{X}\times\Theta\to\mathbb{R}$ are also continuously differentiable w.r.t. the parameter $\theta$ [25], [39]. Since continuous differentiability on the compact set $\Theta\times\mathcal{X}$ implies Lipschitz continuity, these CBFs are Lipschitz continuous. In addition, many CBFs used in safe-control applications are known to be Lipschitz continuous [2], [25], [39], [45].
Definition 2 (Robust CBF).
Let Assumption 1 hold. The continuous barrier function $B_{\theta^{\ast}}:=B(\cdot,\theta^{\ast})$ is a robust DT-CBF for (6) on $\mathcal{S}_{B,\theta^{\ast}}$ if there exists $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$ such that $\forall x\in\mathcal{X}$, $\exists u\in\mathcal{U}$ satisfying
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B_{\theta^{\ast}}(f(x,u;\theta^{\ast}))\hskip-2.27626pt-\hskip-2.27626ptB_{\theta^{\ast}}(x)}\hskip-2.27626pt-\hskip-2.27626ptL_{B,x}\overline{w}\geq-\alpha\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B_{\theta^{\ast}}(x)}\right),\vskip-5.69046pt$ (8)
and (8) is called the robust CBC against disturbances.
The barrier certificate (8) also defines the set of robust safe control inputs for $\mathcal{S}_{B,\theta^{\ast}}$ as
[…]
where the subscript $\theta^{\ast}$ emphasizes that characterizing $\mathcal{U}_{\mathrm{rs},\theta^{\ast}}(x)$ requires knowing the true parameter. The following theorem is a standard result on CBF-based safety guarantees, and the proof is omitted for brevity since it follows the same procedure as in [2].
Theorem 1 (Robust safety).
If the continuous function $B_{\theta^{\ast}}:\mathcal{X}\to\mathbb{R}$ is a robust DT-CBF for (6) on $\mathcal{S}_{B,\theta^{\ast}}$, then $\mathcal{U}_{\mathrm{rs},\theta^{\ast}}(x;\theta^{\ast})$ is non-empty for all $x\in\mathcal{X}$. In addition, if a state-feedback control policy $\pi:\mathcal{X}\to\mathbb{R}^{m}$ satisfies $\pi(x)\in\mathcal{U}_{\mathrm{rs},\theta^{\ast}}(x;\theta^{\ast})$, $\forall x\in\mathcal{X}$, then the closed-loop system $x_{t+1}=f(x_{t},\pi(x_{t});\theta^{\ast})+w_{t}$ is robustly safe w.r.t. $\mathcal{S}_{B,\theta^{\ast}}$ against disturbances $w_{t}$.
Since knowing $\theta^{\ast}$ is generally a stringent requirement, the certificate (8) cannot be directly used to synthesize safe controllers when $\theta^{\ast}$ is unknown. In such cases, a straightforward remedy is to robustify against the worst-case model mismatch and to construct a new inequality condition using a constant nominal $\hat{\theta}$ with additional robustification terms that scale with the estimation error. Accordingly, one can establish a worst-case robust safety guarantee, similar to Theorem 1. However, such worst-case design may react more slowly to dangerous system behaviors compared to those that have access to the true model, and thereby require more conservative control inputs to maintain safety. Since parametric uncertainty is epistemic in nature [40] and can be reduced or even eliminated through online data acquisition, it is desirable to incorporate adaptive safety mechanisms with online model learning to reduce conservatism.
3 Robust adaptive safe control
This section introduces robust adaptive safe control using DT-CBFs with general online parameter estimators that can provide (i) point estimates, (ii) set estimates, and (iii) a uniform bound on the point-estimate increments. More specifically, a sufficient condition used to certify robust adaptive DT-CBFs is provided, and an online input condition is derived to compute safe control inputs in the safety filter framework [12]. The proposed approach can be combined with any nominal controller, enabling modular safe control design.
3.1 Online parameter estimation
Consider a given parameter set $\Theta$, which in practice is constructed from physical bounds or measurement data [1], typically as a box or polytope enclosing all feasible parameter values. At each time step $t$, consider a parameter estimator $\mathtt{EST}$ that can provide a polytopic set estimate $\Theta_{t}\subseteq\Theta$ such that $\theta^{\ast}\in\Theta_{t}$ and a point estimate $\hat{\theta}_{t}\in\Theta_{t}$. The set $\Theta_{t}$ can be obtained, e.g., through set-membership identification, where the feasible parameter set is updated recursively by intersecting it with the sets consistent with new measurements at each time step. The initial set $\Theta_{0}$ is usually taken as $\Theta$. Following (4) and (5), the instantaneous estimation error and its associated error bound are defined as follows:
$e_{\theta,t}:=\hat{\theta}_{t}-\theta^{\ast},$ (9a)
$\varepsilon_{\theta,t}(p):=\max_{\theta\in\Theta_{t}}\|\theta-\hat{\theta}_{t}\|_{p}.\vskip-8.5359pt$ (9b)
The module $\mathtt{EST}$ performs a recursive point-estimate update specified by the following abstracted update law:
$\hat{\theta}_{t+1}=\mathcal{T}_{\mathtt{ADP},t}(\hat{\theta}_{t}),\vskip-5.69046pt$ (10)
where $\mathcal{T}_{\mathtt{ADP},t}$ is time-varying since the closed-loop input-state data is collected online and the set estimate $\Theta_{t}$ will also be used to determine the point estimate. Finally, the point-estimate increment is defined by
$\delta_{\theta,t}:=\hat{\theta}_{t+1}-\hat{\theta}_{t}=\mathcal{T}_{\mathtt{ADP},t}(\hat{\theta}_{t})-\hat{\theta}_{t}.\vskip-5.69046pt$ (11)
Moreover, the considered estimator $\mathtt{EST}$ guarantees an uniform upper bound on the increments, i.e.,
$\|\delta_{\theta,t}\|\leq\bar{\delta}_{\theta}.\vskip-5.69046pt$ (12)
Note that such a bound always exists since $\Theta$ is polytopic and therefore compact, implying that $\|\delta_{\theta,t}\|$ always admits a worst-case bound given by $\max_{\theta^{\prime},\theta^{\prime\prime}\in\Theta}\|\theta^{\prime}-\theta^{\prime\prime}\|$.
3.2 Robust adaptive safe control
$\mathcal{X}$$\mathcal{S}_{\text{ra},t}$$\mathcal{S}_{B,\hat{\theta}_{t}}$
(a) Robustification
$\pi_{\text{nom},t}(x)$$\pi_{\text{ras},t}(x;p)$$\mathcal{U}_{\text{ras},t}(x;p)$$\mathcal{U}$
(b) Safety filter
Figure 1: Sketch of (a) adaptively robustified safe sets, and (b) the safety filter projection when the nominal input is unsafe.
Given a CBF $B:\mathcal{X}\times\Theta\to\mathbb{R}$ that encodes safety constraints and that defines a safe set as in (7), we define its corresponding adaptively robustified counterpart by
$B_{\mathrm{ra},t}(x):={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}B(x,\hat{\theta}_{t})}-\frac{1}{2}e_{\theta,t}^{\top}\Gamma^{-1}e_{\theta,t},\vskip-5.69046pt$ (13)
where $\Gamma$ is a symmetric positive definite matrix. Accordingly, the robust adaptive safe set $\mathcal{S}_{\mathrm{ra},t}$ is defined by the 0-superlevel set of $B_{\mathrm{ra},t}(\cdot)$ as
$\mathcal{S}_{\mathrm{ra},t}:=\{x\in\mathcal{X}\mid B_{\mathrm{ra},t}(x)\geq 0\}.\vskip-5.69046pt$ (14)
It is noted that $\mathcal{S}_{\mathrm{ra},t}\subseteq\mathcal{S}_{B,\hat{\theta}_{t}}$ (see Fig. 1(a)), and $\mathcal{S}_{\mathrm{ra},t}=\mathcal{S}_{B,\theta^{\ast}}$ if and only if $\hat{\theta}_{t}=\theta^{\ast}$. Since $\mathcal{S}_{\mathrm{ra},t}$ is time-varying, a timed version of Definition 1 is given as follows:
Definition 3.1 (Timed robust safety).
Given $T\in\overline{\mathbb{N}}$, $\{\mathcal{X}_{\mathrm{s},t}\}^{T}_{t=0}$ satisfying $x_{0}\in\mathcal{X}_{\mathrm{s},0}$ and $\forall t\in\mathbb{I}_{[0:T]}$, $\mathcal{X}_{\mathrm{s},t}\subseteq\mathcal{X}$ is sequentially positively invariant for (6) if $\forall t\in\mathbb{I}_{[0:T-1]}$, $x_{t}\in\mathcal{X}_{\mathrm{s},t}$ implies that $x_{t+1}\in\mathcal{X}_{\mathrm{s},t+1}$ for all $w_{t}\in\mathcal{W}$. In this case, the system (6) is said to be robustly safe w.r.t. $\{\mathcal{X}_{\mathrm{s},t}\}^{T}_{t=0}$. In addition, if $\mathcal{X}_{\mathrm{s},t}\subseteq\mathcal{X}_{\mathrm{s}}^{\ast}$, $\forall t\in\mathbb{I}_{[0:T]}$, the system (6) is said to be robustly safe w.r.t. $\mathcal{X}_{\mathrm{s}}^{\ast}$ over $\mathbb{I}_{[0:T]}$.
The concept of imposing finite-horizon safety has been used in safe control using CBFs [10], [11]. In fact, the timed safety in Definition 3.1 is more practical and general than its infinite-horizon counterpart since safety usually needs only be ensured over the duration of operation [34] and taking $T=+\infty$ recovers infinite-horizon safety.
Since $\theta^{\ast}$ is unknown, $e_{\theta,t}$ cannot be computed, and thus $\mathcal{S}_{\mathrm{ra},t}$ cannot be explicitly characterized. However, $\mathcal{S}_{\mathrm{ra},t}$ only serves as a fictitious set to realize safety characterized by $B$, as will be manifested in Theorem 2. Next, we present the robust adaptive CBC, which can be used to certify robust adaptive DT-CBFs. Moreover, given such a valid DT-CBF, safe control inputs can be computed in the safety filter framework [3] for systems subject to both bounded disturbances and parametric uncertainty.
Definition 3.2 (Robust adaptive CBF).
Let Assumption 1 hold. Given an estimator $\mathtt{EST}$ and $T\in\overline{\mathbb{N}}$, a continuous function $B:\mathcal{X}\times\Theta\to\mathbb{R}$ is a robust adaptive DT-CBF for (6) and $\mathtt{EST}$ if there exists $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$ such that, for all $x\in\mathcal{X}$, there exists $u\in\mathcal{U}$ satisfying
$\Delta B{{}^{\star}}(x,u)-L_{B,x}\overline{w}-E^{\star}_{\theta}(x)\\ \geq-\alpha\left(B^{\star}(x)-\frac{(\bar{\varepsilon}_{\theta}(p))^{2}}{2\lambda_{\min}(\Gamma)}\right)\vskip-5.69046pt$ (15)
where $\Gamma$ is given as in (13), $B^{\star}(x):=\min_{\theta\in\Theta}B(x,\theta)$, $\Delta B{{}^{\star}}(x,u)\hskip-2.27626pt:=\hskip-2.27626pt\min_{\theta\in\Theta}\{B(f(x,u;\theta),\theta)\hskip-2.27626pt-\hskip-2.27626ptB(x,\theta)\}$, $\bar{\varepsilon}_{\theta}(p)$ is the worst-case error $p$-norm ($p\in\{1,2\}$) defined as
$\bar{\varepsilon}_{\theta}(p):=\max_{\theta,\theta^{\prime}\in\Theta}\|\theta-\theta^{\prime}\|_{p},\vskip-5.69046pt$ (16)
and $E^{\star}_{\theta}(x)=\left(L_{B,x}\|\bm{\phi}(x)\|+\frac{\bar{\delta}_{\theta}}{\lambda_{\min}(\Gamma)}\right)\bar{\varepsilon}_{\theta}(p)+\frac{\bar{\delta}_{\theta}^{2}}{2\lambda_{\min}(\Gamma)}+L_{B,\theta}\bar{\delta}_{\theta}$, with $L_{B,x}$ and $L_{B,\theta}$ given in Assumption 1 and $\bar{\delta}_{\theta}$ characterized by $\mathtt{EST}$ as in (12). In this case, the condition (15) is called the robust adaptive CBC against disturbances and parametric model mismatch.
Theorem 2 (Robust adaptive safety).
Let Assumption 1 hold and assume $x_{0}\in\mathcal{S}_{\mathrm{ra},0}$. Given an estimator $\mathtt{EST}$ and $T\in\overline{\mathbb{N}}$, if $B$ is a robust adaptive DT-CBF for (6) and $\mathtt{EST}$ with its associated $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$, then for the robust adaptive online input condition
$B\big(f(x,u;\hat{\theta}_{t}),\hat{\theta}_{t}\big)-B(x,\hat{\theta}_{t})-L_{B,x}\overline{w}\\ -E_{\theta,t}(x)\geq-\alpha\left(B(x,\hat{\theta}_{t})-\frac{(\varepsilon_{\theta,t}(p))^{2}}{2\lambda_{\min}(\Gamma)}\right),\vskip-11.38092pt$ (17)
the set $\mathcal{U}_{\mathrm{ras},t}(x;p)$ defined as
[…]
is non-empty, $\forall x\in\mathcal{X}$ and $t\in\mathbb{I}_{[0,T-1]}$. The term $E_{\theta,t}(x)$ in (17) equals $\left(L_{B,x}\|\bm{\phi}(x)\|+\frac{\|\delta_{\theta,t}\|}{\lambda_{\min}(\Gamma)}\right)\varepsilon_{\theta,t}(p)$ $\frac{\|\delta_{\theta,t}\|^{2}}{2\lambda_{\min}(\Gamma)} +$$L_{B,\theta}\|\delta_{\theta,t}\|$ with $\delta_{\theta,t}$ defined in (11). In addition, if time-varying state-feedback control policies $\pi_{t}:\mathcal{X}\to\mathbb{R}^{m},t\in\mathbb{I}_{[0,T-1]}$ satisfies $\pi_{t}(x)\in\mathcal{U}_{\mathrm{ras},t}(x;p)$ for all $x\in\mathcal{X}$, then the time-varying closed-loop system
$x_{t+1}=f(x_{t},\pi_{t}(x_{t});\theta^{\ast})+w_{t}\vskip-5.69046pt$ (19)
is robustly safe w.r.t. $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$ and thus robustly safe w.r.t. $\{\mathcal{S}_{B,\hat{\theta}_{t}}\}^{T}_{t=0}$ (cf. Definition 3.1) against both parametric model mismatch and disturbances.
Verifying a candidate barrier function using the certificate (15) does not require the knowledge of $\theta^{\ast}$. Note that (15) is a discrete-time counterpart of the certificate in [25] for continuous-time systems. The proof of Theorem 2 is given in Appendix A. It is worth noting that the online safe control condition (17) requires only the estimated parameters without needing knowledge of $\theta^{\ast}$. This property makes (17) readily applicable for computing safe inputs via a safety filter, one that evaluates each nominal input and replaces it with the closest admissible input that renders the system safe (see Fig. 1(b)). Specifically, given a nominal state-feedback policy $\pi_{\mathrm{nom},t}:\mathcal{X}\to\mathbb{R}^{m}$ at time step $t$, a robust adaptive safe policy $\pi_{\mathrm{ras},t}:\mathcal{X}\to\mathcal{U}$ is implicitly given by:
$\pi_{\mathrm{ras},t}(x;p)=\operatorname*{arg\,min}_{\nu\in\mathcal{U}_{\mathrm{ras},t}(x,p)}\|\nu-\pi_{\mathrm{nom},t}(x)\|.\vskip-5.69046pt$ (20)
The safe policy $\pi_{\mathrm{ras},t}(\cdot;p)$ is time-varying partly because the underlying admissible set $\mathcal{U}_{\mathrm{ras},t}$ evolves over time, inherently capturing the influence of past states and inputs via the estimator $\mathtt{EST}$.
Remark 3.3.
The online input condition (17) is less conservative than the CBC (15), which is only used for offline design and verification of CBFs. Besides, the robustified barrier function (13) is built using the true parametric error $e_{\theta,t}$ instead of the set-induced error bound as in [25]. In fact, unlike what has been reported in [25], the invariance property of the safe sets defined by the maximized error bound-based robust adaptive CBFs can only be derived in discrete time, but not in continuous time. More details are given in Appendix B.
3.3 Practical implementation
Deriving a closed-form $\pi_{\mathrm{ras},t}(\cdot;p)$ is difficult in general. In practice, it suffices to compute $u_{\mathrm{ras},t}(p)=\pi_{\mathrm{ras},t}(x_{t};p)$ by solving (20) for the current state $x_{t}$. This relies on characterizing $\mathcal{U}_{\mathrm{ras},t}(x_{t};p)$ via the condition (15), which also depends on $\varepsilon_{\theta,t}(p)$. Computing $\varepsilon_{\theta,t}(2)$ generally requires solving a non-convex program, while using $\varepsilon_{\theta,t}(1)$ reduces (9b) to a linear program that can be solved efficiently. Since $\varepsilon_{\theta,t}(2)\leq\varepsilon_{\theta,t}(1)$, we have $\mathcal{U}_{\mathrm{ras},t}(x;1)\subseteq\mathcal{U}_{\mathrm{ras},t}(x;2)$, meaning the $1$-norm-based condition is more conservative but offers faster computation in practice.
In general, the optimization problem (20) is non-convex and cannot be solved efficiently. However, noting that $f$ is linear in $u$ due to the input-affine structure (cf. the model in (2.2)), when the designed barrier candidate $B$ is concave w.r.t. its first argument such that $\mathcal{U}_{\mathrm{ras},t}(x;p)$ is a convex set, then computing $u_{\mathrm{ras},t}(p)$ is equivalent to solving a convex optimization problem, for which many efficient algorithms with convergence and optimality guarantees exist [7]. Note that the concavity of $B$ w.r.t. the state $x$ is a standard assumption in many DT-CBF formulations [2], [44]; however, the design of concave model-based barrier functions depends on specific applications, which is out of the scope of the current paper.
Remark 3.4.
Employing $\varepsilon_{\theta,t}(1)$ in (17) results in a more conservative safe controller. Thus, $u_{\mathrm{ras},t}(1)$ is likely to be more suboptimal than $u_{\mathrm{ras},t}(2)$, especially in the context of optimal control when the policy $\pi_{\mathrm{nom},t}$ is (implicitly) given by methods using discrete-time control Lyapunov functions (CLFs) [5], reinforcement learning [9], or approximate MPC [13], [20]. An alternative approach is to use the $\infty$-norm uniformly when quantifying all the error bounds (i.e., (5), (9b), (12), and (16)) and the Lipschitz condition in Assumption 1.
4 Online estimation & Causal Synthesis
In this section, following the adaptive control approach for linear systems [26], we instantiate an estimator $\mathtt{EST}$ using the clipped recursive-least-square (RLS) update with set-membership projection. In addition, we present a causal integration of the safety filter and the estimator to guarantee that computing $u_{\mathrm{ras},t}(p)$ does not rely on any information from time step $t+1$ onward.
For $t\geq 1$, both $\hat{\theta}_{t}$ and $\Theta_{t}$ are determined by the information flow $\mathcal{F}_{t-1}:=(x_{0},u_{0},x_{1},u_{1},\dots,x_{t-1})$, and a causal synthesis means computing $u_{t}$ based on $\mathcal{F}_{t}$. The initial estimate satisfies $\hat{\theta}_{0}\in\Theta_{0}$ with $\Theta_{0}=\Theta$. Given $x_{t}$, the residual is defined by $r_{t}:=x_{t}-f_{\mathrm{d}}(x_{t-1})-g(x_{t-1})u_{t-1}$, and it holds that $r_{t}+[\bm{\phi}(x_{t-1})]^{\top}\theta^{\ast}\in\mathcal{W}$, i.e., $H_{w}(r_{t}+[\bm{\phi}(x_{t-1})]^{\top}\theta^{\ast})\leq\mathbf{1}_{n_{w}}$ by (2). Equivalently, $\theta^{\ast}\in\Delta_{\theta,t}$ with the non-falsified polytopic set $\Delta_{\theta,t}$ given by
$\Delta_{\theta,t}=\left\{\theta\in\mathbb{R}^{q}\mid H_{w}[\bm{\phi}(x_{t-1})]^{\top}\theta\leq\mathbf{1}_{n_{w}}\hskip-2.27626pt-\hskip-2.27626ptH_{w}r_{t}\right\}.\vskip-5.69046pt$ (21)
Given the current set estimate $\Theta_{t}=\{\theta\in\mathbb{R}^{q}\mid H_{\theta,t}\theta\leq h_{\theta,t}\}$, the updated set estimate is computed as
$\Theta_{t+1}=\Theta_{t}\cap\Delta_{\theta,t}\vskip-5.69046pt$ (22)
which is a polytope described by an augmented set of hyperplanes [^5] Unbounded growth in the number of hyperplanes can be avoided by overapproximating the intersection with a hyperrectangle (or hyberbox) of fixed dimension [26].contributing from $\Theta_{t}$ and $\Delta_{\theta,t}$. For the point estimate, a prior is first computed based on the clipped RLS parameter update as
$\hat{\theta}_{t+1}^{[\mathrm{p}]}=\hat{\theta}_{t}-\mu_{t}\underbrace{\bm{\phi}(x_{t-1})\left(x_{t}-f(x_{t-1},u_{t-1};\hat{\theta}_{t})\right)}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}:=\psi_{\theta,t}}},\vskip-5.69046pt$ (23)
where $\mu_{t}=\min\left\{(\left\|\psi_{\theta,t}\right\|+\epsilon)^{-1},\|\bm{\phi}(x_{t-1})\|^{-2}\right\}$ is the clipped learning rate with $\epsilon>0$ being a small constant. Note that (23) degenerates to a gradient descent update when the learning rate is constant [26], and the proposed clipped RLS update follows gradient clipping [16]. The final updated estimate $\hat{\theta}_{t+1}$ is computed by set-membership projection as
$\hat{\theta}_{t+1}=\Pi_{\Theta_{t+1}}(\hat{\theta}_{t+1}^{[\mathrm{p}]}),\vskip-5.69046pt$ (24)
and $\delta_{\theta,t}$ in (11) can be obtained accordingly. Due to $\|\mu_{t}\psi_{\theta,t}\|\leq 1$ and the triangle inequality, it holds that $\|\delta_{\theta,t}\|\leq 1$. Since $\delta_{\theta,t}$ is determined by $\mathcal{F}_{t}$, and thus $u_{t}$, as computed as in (20), also depends on $\mathcal{F}_{t}$. In short, causality is ensured by synchronizing the estimates $\hat{\theta}_{t}$ and $\Theta_{t}$ with the one-step-delayed information flow $\mathcal{F}_{t-1}$.
Algorithm 1 outlines the core steps of the proposed safe control approach. The required inputs are the CBF $B$ and its associated $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$, initial variables $x_{0}$, $\hat{\theta}_{0}$ and $\Theta$, and constants $\overline{w}$, $L_{B,x}$, $L_{B,\theta}$, and $\Gamma$. At time step $t$, $\hat{\theta}_{t}$ and $\delta_{\theta,t}$ are first computed to define $\mathcal{U}_{\mathrm{ras},t}(x_{t};p)$, and $\hat{\theta}_{t+1}$ and $\Theta_{t+1}$ will be used at $t+1$. Given $\mathcal{U}_{\mathrm{ras},t}(x;p)$, a safe $u_{t}$ is then computed using the safety filter, and the system state evolves to $x_{t+1}$. It is also worth noting that $\Theta_{1}=\Theta_{0}$ and $\hat{\theta}_{1}=\hat{\theta}_{0}$, meaning the non-trivial update starts from $t=1$.
Algorithm 1 Robust Adaptive Safe Control
$B$, $\alpha$, $x_{0}$, $\hat{\theta}_{0}$, $\Theta$, $\overline{w}$, $L_{B,x}$, $L_{B,\theta}$, $\Gamma$
$\Theta_{0}\leftarrow\Theta$, $\mu_{0}\leftarrow 0$, $\Delta_{\theta,0}\leftarrow\mathbb{R}^{p}$
for $t=0$ to T1T-1 do
  Obtain nominal policy $\pi_{\mathrm{nom},t}$
  Compute $\varepsilon_{\theta,t}(p)$ $\triangleright$ (9b)
  Compute $\hat{\theta}_{t+1}$ and $\delta_{\theta,t}$ $\triangleright$ (21)–(24), (11)
  Determine $\mathcal{U}_{\mathrm{ras},t}(x;p)$ $\triangleright$ (15)
  Compute $u_{t}$ using the safety filter $\triangleright$ (20)
  Obtain $x_{t+1}$ by applying $u_{t}$
end for
Remark 4.1.
The nominal policies $\{\pi_{\mathrm{nom},t}\}^{T-1}_{t=0}$ can also be designed by combining the estimator $\mathtt{EST}$ if model-based approaches are adopted, following the certainty-equivalence principle [21], [22] in adaptive control with online estimation [23], [26].
Remark 4.2.
Convergence of the estimate $\hat{\theta}_{t}$ to the true parameter $\theta^{\ast}$ typically requires additional conditions such as persistent excitation [24], [29], which can be achieved through active exploration (e.g., injecting noise into the input [9]). In contrast, the proposed safe control approach relies only on the uncertainty set $\Theta_{t}$ that contains the true parameter $\theta^{\ast}$, such that safety is ensured at all times by robustification against $\Theta_{t}$. Although convergence allows conservatism to be reduced maximally as $E_{\theta,t}(x)$ in (17) approaches zero, the safety guarantees developed in this work do not rely on achieving convergence. The explicit design of exploration strategies to ensure convergence is therefore beyond the scope of the current paper.
5 Inherent robustness
Inspired by [43], we discuss the inherent robustness property of the input condition in (17). Specifically, when $x_{t_{0}}$ is only safe but not robustly safe (i.e., $x_{t_{0}}\in\mathcal{S}\setminus\mathcal{S}_{\mathrm{ra},t_{0}}$) or even unsafe (i.e., $x_{t_{0}}\notin\mathcal{S}$) for some $t_{0}\in\mathbb{N}$ due to unexpected perturbations [43] or poor parameter estimates, inherent robustness ensures that the safe inputs derived from the safety filter (20) based on the condition (17) are able to steer the state towards the sequence of robust adaptive safe sets $\{\mathcal{S}_{\mathrm{ra},t}\}^{\infty}_{t=t_{0}}$ asymptotically, i.e., $x_{t}\in\mathcal{S}_{\mathrm{ra},t}$ when $t\to\infty$. This asymptotic convergence w.r.t. a sequence of sets is formally defined as follows:
Definition 5.1.
Assume that $\{\mathcal{X}_{t}\}^{\infty}_{t=0}$ with $\mathcal{X}_{t}\subseteq\overline{\mathcal{X}}$ is sequentially positively invariant (cf. Definition 3.1) for a general time-varying system $x_{t+1}=h_{t}(x_{t})+w_{t}$ for all $w_{t}\in\mathcal{W}$, where $h_{t}:\mathcal{X}\to\mathbb{R}^{n}$ is continuous for all $t\in\overline{\mathbb{N}}$. Then $x_{t}$ converges to $\{\mathcal{X}_{t}\}^{\infty}_{t=0}$ asymptotically on $\overline{\mathcal{X}}$ is defined as $\lim_{t\to\infty}\mathrm{dist}_{\mathcal{X}_{t}}(x_{t})=0$ for all $x_{0}\in\overline{\mathcal{X}}$.
Following (14), we define the corresponding time-varying continuous energy functions on $\mathcal{X}$ as follows:
$V_{t}(x)=\begin{cases}0&\text{if }x\in\mathcal{S}_{\mathrm{ra},t}\\ -B_{\mathrm{ra},t}(x)&\text{if }x\in\mathcal{X}\setminus\mathcal{S}_{\mathrm{ra},t}\end{cases},\quad\forall t\in\overline{\mathbb{N}}.\vskip-5.69046pt$ (25)
It is obvious that $\forall t\in\overline{\mathbb{N}},V_{t}(x)>0$ for $x\in\mathcal{X}\setminus\mathcal{S}_{\mathrm{ra},t}$. Given that $\mathcal{S}_{\mathrm{ra},t}$ is sequentially positively invariant (see Appendix A), $V_{t}$ can be used to show that $x_{t}$ converges to $\{\mathcal{S}_{\mathrm{ra},t}\}^{\infty}_{t=0}$ asymptotically if the input $u_{t}$ satisfies (15), which is stated in the following theorem.
Theorem 3 (Inherent robustness guarantee).
Let Assumption 1 hold. If $B:\mathcal{X}\times\Theta\to\mathbb{R}$ is a robust adaptive DT-CBF for (6) and $\mathtt{EST}$ (cf. Definition 3.2), then $x_{t}$ converges to $\{\mathcal{S}_{\mathrm{ra},t}\}^{\infty}_{t=0}$ asymptotically on $\mathcal{X}$ (cf. Definition 5.1) for the system (6) with $u_{t}\in\mathcal{U}_{\mathrm{ras},t}(x_{t};p)$.
The proof of Theorem 3 is provided in Appendix C. Theorem 3 directly implies that a sequence of inputs certified by (15) ensure that $x_{t}\in\mathcal{S}_{\mathrm{ra},t}$ when $t\to\infty$. Moreover, although $\mathcal{S}_{\mathrm{ra},t}$ does not generally converge to $\mathcal{S}$, the asymptotic convergence of $\{\mathcal{S}_{\mathrm{ra},t}\}^{\infty}_{t=0}$ is not affected. As discussed in Remark 3.3, maximized error bound-based CBFs [25] can also be used to establish safety, though their induced CBC is more conservative than (15). However, the maximized error bound can exhibit a monotonicity property [25], which can cause the resulting sequence of safe sets to converge to $\mathcal{S}$ under mild conditions. In this case, the inherent robustness property also holds, and more details are given in Appendix C.
6 Application Example
[…]
Figure 2: Diagram of the proposed safe control approach for the adaptive cruise control problem using robust adaptive discrete-time control barrier functions.
In this section, we apply the proposed DT-CBF-based safe control to an adaptive cruise control (ACC) problem, a standard benchmark in safe control [3], [39], [42]. The ego vehicle dynamics are
[…]
where $v$, $M$, and $d$ denote velocity, vehicle mass, and the distance to the front vehicle, respectively. The nominal velocity of the front vehicle $v_{\mathrm{f}}\in[v_{\text{lb}},v_{\text{ub}}]$ is unknown. The input $u\in[-u_{\max},u_{\max}]$ represents the traction force applied by the ego vehicle. The resistance force $F_{\mathrm{r}}(v)$ accounts for drag effects and is modeled as $F_{\mathrm{r}}(v)=F_{\text{roll}}+\mu_{\text{vis}}v+\mu_{\text{aero}}v^{2}$ for $v>0$, where $F_{\text{roll}}$ is the Coloumb friction force, $\mu_{\text{vis}}$ is the viscous force coefficient, and $\mu_{\text{aero}}$ is the aerodynamic drag coefficient [39], [42]. Both $F_{\text{roll}}$ and $\mu_{\text{vis}}$ are determined empirically offline [42], whereas $\mu_{\text{aero}}\in[\mu^{-},\mu^{+}]$ is identified online. The disturbance $w_{v}\in[-\overline{w}_{v},\overline{w}_{v}]$ is included to capture wind effects, and $w_{d}\in[-\overline{w}_{d},\overline{w}_{d}]$ characterizes variations of $v_{\mathrm{f}}$. Table 1 summarizes the nominal parameters and the known bounds. For simplicity and computational efficiency [11], [38], Euler discretization with sampling period $\Delta t=0.1[$\mathrm{s}$\$] is applied to obtain
[…]
The model (27) is of the general form in (2.2) with box-constrained $w_{t}=[w_{v,t}\Delta t,w_{d,t}\Delta t]^{\top}$ and $\theta=[\mu_{\text{aero}},v_{\mathrm{f}}]^{\top}$ satisfying (2) and (3). Safety is enforced through the time-headway constraint $d\geq 1.8v$ [39], and the CBF [^6] For the considered adaptive cruise control example, we adopt the parameter-free CBF as in [39].is chosen as B(v,d)=d1.8v0.5B(v,d)=d-1.8v-0.5 with $\alpha(x)=(1-10^{-4})x$. It can be numerically verified that the CBC in (15) is satisfied for $p=2$ and $\Gamma=10^{5}\mathbf{I}_{2}$. The nominal controller is an unconstrained myopic (i.e., prediction horizon $N=1$) adaptive MPC (aMPC) controller [^7] Our contribution is not centered on MPC; it is used only as a nominal controller for numerical illustration. tracking $v_{\text{ref}}=30[\text{m/s}]$, equiped with the estimator that computes $\hat{\mu}_{\text{aero}}$ and $\hat{v}_{\text{f}}$ as in Section 4. The nominal input from aMPC is filtered by the safety filter in (20) to obtain the safe input. Fig. 2 presents a diagram that highlights the connections between the theoretical components and their implementation.
Figure 3: $\mathtt{EST}$ performance. The shaded area is the min-max envelope encompassing different disturbance realizations.
Figure 4: Comparison of adaptive MPC with robust adaptive CBF-based safety filter (aMPC-raCBF), MPC with worst-case robust CBF-based safety filter (MPC-rCBF), adaptive MPC (aMPC), and robust adaptive MPC (raMPC) [18]. The shaded area is the min-max envelope encompassing different disturbance realizations.
Table 1: Parameters and bounds used in the numerical simulation.
Nominal parameters Bounds
Parameter Value Units Parameter Value Units
$M$ $1650$ $\mathrm{kg}$ $v_{\text{lb}},v_{\text{ub}}] [$2020, $32$] $\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$
$F_{\text{roll}}$ $125$ $\mathrm{kg}\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$ $\mu^{-},\mu^{+}] [$$0.1$, $0.6$] $\mathrm{kg}\text{\,}{\mathrm{m}}^{-1}$
$\mu_{\text{vis}}$ $1.2$ $\mathrm{kg}\text{\,}{\mathrm{s}}^{-1}$ $\overline{w}_{v}$ 22 $\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$
$\mu^{\ast}_{\text{aero}}$ $0.55$ $\mathrm{kg}\text{\,}{\mathrm{m}}^{-1}$ $\overline{w}_{d}$ 55 $\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$
$v^{\ast}_{\text{f}}$ $22$ $\mathrm{m}\text{\,}{\mathrm{s}}^{-1}$ $u_{\max}$ 10 00010\,000 $\mathrm{kg}\text{\,}\mathrm{m}\text{\,}{\mathrm{s}}^{-2}$
The simulations [^8] The full codebase and specific implementations are available at: https://github.com/lcrekko/ra_cbf_mpc. are conducted using Python 3.12. The performance of the online estimator $\mathtt{EST}$ is illustrated in Fig. 3, which demonstrates that the true parameters are accurately recovered over time. The results related to safety are presented in Fig. 4. As shown in the middle subplot, the proposed robust adaptive CBF-based safety filter ensures constraint satisfaction even in the presence of disturbances and model mismatch. This is achieved by actively modifying unsafe control inputs when safety is at risk, as evidenced in the bottom subplot of the same figure. Based on Fig. 4, the need for an explicit safety mechanism becomes apparent when comparing with the baseline unconstrained aMPC controller, which fails to enforce the safety constraint. Although method using robust CBF and robust adaptive MPC with constraint tightening [18] also maintains safety, they react more slowly to critical situations compared to the safe control approach using robust adaptive CBFs. As a result, from approximately $7$ seconds onward, stronger braking actions–i.e., more conservative control inputs–are required to uphold safety. Although robust adaptive MPC [17], [18], [27] provides an alternative approach to safe control under uncertainty, methods based on CBFs are modular by design and can be combined with nominal controllers that do not inherently guarantee safety, e.g., reinforcement learning [8] and approximate MPC [13], [20].
7 Conclusions & Future work
In this paper, we have developed an adaptive safe control method for discrete-time systems using control barrier functions, considering both parametric model uncertainty and disturbances. Unlike traditional continuous-time adaptive safe control designs that tightly couple adaptation and control, our framework supports a modular architecture such that the parameter estimator operates independently of the safety filter. Future research directions include developing safe control methods for output-feedback systems with imperfect state measurements, extending the framework to systems with nonlinearly parameterized uncertainty, incorporating active exploration mechanisms for safe learning under model uncertainty, and characterizing the existence of CBFs and designing CBFs for uncertain systems.
Appendix A Proof of Theorem 2
{pf}
First, due to the definition of $\Delta B^{\star}(x,u)$ and $B^{\star}(x)$, the increment bound (12), and $\bar{\varepsilon}_{\theta}(p)\geq\varepsilon_{\theta,t}(p)$, it can be obtained that (15) is a sufficient condition for (17). Therefore, if $B$ is a robust adaptive DT-CBF that satisfies (15), then $\mathcal{U}_{\mathrm{ras},t}(x;p)$ is non-empty. Next, it suffices to prove robust safety through positive invariance. Considering $B_{\mathrm{ra},t+1}(x_{t+1})-B_{\mathrm{ra},t}(x_{t})$, which characterizes the increment of the barrier function, it holds that
$B_{\mathrm{ra},t+1}(x_{t+1})-B_{\mathrm{ra},t}(x_{t})$
$=$ $B(x_{t+1},\hat{\theta}_{t+1})-B(x_{t},\hat{\theta}_{t})$
$-\frac{1}{2}(e_{\theta,t+1}-e_{\theta,t})^{\top}\Gamma^{-1}(e_{\theta,t+1}+e_{\theta,t})$
[…]
$\geq$ $B(x_{t+1},\hat{\theta}_{t+1})-B(x_{t},\hat{\theta}_{t})-\frac{\|\delta_{\theta,t}\|^{2}+2\|\delta_{\theta,t}\|\varepsilon_{\theta,t}(p)}{2\lambda_{\min}(\Gamma)},\vskip-14.22636pt$ (28)
where $B(x_{t+1},\hat{\theta}_{t+1})-B(x_{t},\hat{\theta}_{t})$ can be further relaxed using the Lipschitz perturbation bound as
$B(x_{t+1},\hat{\theta}_{t+1})-B(x_{t},\hat{\theta}_{t})$
$\hskip-12.91663pt\geq B(f(x_{t},\pi_{t}(x_{t});\hat{\theta}_{t}),\hat{\theta}_{t})-B(x_{t},\hat{\theta}_{t})$
$-L_{B,x}(\|\bm{\phi}(x_{t})\|\varepsilon_{\theta,t}(p)+\overline{w})-L_{B,\theta}\|\delta_{\theta,t}\|.\vskip-5.69046pt$ (29)
By substituting (A) into (A) and using the fact that $\pi_{t}(x_{t})$ satisfies (15), we have
$\hskip-8.61108ptB_{\mathrm{ra},t+1}(x_{t+1})-B_{\mathrm{ra},t}(x_{t})\\ \hskip-17.22217pt\geq-\alpha\left(B(x_{t},\hat{\theta}_{t})\hskip-2.27626pt-\hskip-2.27626pt\frac{(\varepsilon_{\theta,t}(p))^{2}}{2\lambda_{\min}(\Gamma)}\right)\geq-\alpha(B_{\mathrm{ra},t}(x_{t})).\vskip-5.69046pt$ (30)
Based on (30), if $x_{t}\in\mathcal{S}_{\mathrm{ra},t}$, i.e., $B_{\mathrm{ra},t}(x_{t})\geq 0$, then $B_{\mathrm{ra},t+1}(x_{t+1})\geq B_{\mathrm{ra},t}(x_{t})-\alpha(B_{\mathrm{ra},t}(x_{t}))\geq 0$ since $\alpha\in\mathcal{C}_{\mathrm{SL}}$, which implies $x_{t+1}\in\mathcal{S}_{\mathrm{ra},t+1}$. Thus, $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$ is sequentially positive invariant for the closed-loop system (19) given $\pi_{t}(x)\in\mathcal{U}_{\mathrm{ras},t}(x;p)$. Moreover, noting that $B_{\mathrm{ra},t}(x)\geq 0\Longrightarrow B(x,\hat{\theta}_{t})\geq 0$, then $\{\mathcal{S}_{B,\hat{\theta}_{t}}\}^{T}_{t=0}$ is positive invariant for (19).               ∎
Appendix B Robustified CBFs with error bounds
The true error $e_{\theta,t}$ is used to construct $B_{\mathrm{ra},t}$ in (13) to achieve adaptive safety, which is similar to [39]. Although safety can be established, $\{\mathcal{S}_{\mathrm{ra},t}\}^{T}_{t=0}$ cannot be characterized since $e_{\theta,t}$ is unknown. Considering that the ultimate goal is to achieve safety w.r.t. $\mathcal{S}_{B,\hat{\theta}_{t}}$, which encodes safety [25], [39], the proposed method using $B_{\mathrm{ra},t}$ defined in (13) is still valid since $\mathcal{S}_{\mathrm{ra},t}\subseteq\mathcal{S}_{B,\hat{\theta}_{t}}$.
An alternative is using the set-induced error bound $\eta_{\theta,t}$ satisfying $|e_{\theta,t}|\leq\eta_{\theta,t}$ to construct robustified CBFs [25]. Specifically, for all $i\in\mathbb{I}_{[1:q]}$, $\eta_{\theta,t}[i]$ can be computed as
$\eta_{\theta,t}[i]=\max_{\theta\in\Theta_{t}}\theta[i]-\min_{\theta\in\Theta_{t}}\theta[i].\vskip-5.69046pt$ (31)
The bound in (31) is conservative since it is solely computed using $\Theta_{t}$. Besides, the $\max$ ($\min$) operation prevents obtaining an analytical expression of its time derivative, and proving set invariance in continuous-time settings has to resort to the true error instead of the bound itself that is used to construct the CBF (see the proof of [25, Theorem 2]). The consequence is that, in the continuous-time setting, it is impossible [^9] Noting that, in the proof of [25, Theorem 2], $h\geq h_{r}-\frac{1}{2}\tilde{\vartheta}^{\top}\Gamma^{-1}\tilde{\vartheta}$ and $h\geq 0$ cannot lead to $h_{r}-\frac{1}{2}\tilde{\vartheta}^{\top}\Gamma^{-1}\tilde{\vartheta}\geq 0$. to establish the invariance property of the safe sets defined by the robustified CBFs constructed using the error bound. Luckily, in the discret-time setting, such invaraince property can be established since no derivative is required. Define the error-bound robustified CBF as
$\overline{B}_{\mathrm{ra},t}(x):=B(x,\hat{\theta}_{t})-\frac{1}{2}\eta_{\theta,t}^{\top}\Gamma^{-1}\eta_{\theta,t},\vskip-5.69046pt$ (32)
where $\Gamma$ is a symmetric positive definite matrix as in (13), and the safe set corresponding to $\overline{B}_{\mathrm{ra},t}$ is given by
$\overline{\mathcal{S}}_{\mathrm{ra},t}=\{x\in\mathcal{X}\mid\overline{B}_{\mathrm{ra},t}(x)\geq 0\}.\vskip-5.69046pt$ (33)
Following Definition 3.2, a modified version based on (32) and (33) is given as follows:
Definition B.1.
Let Assumption 1 hold. Given an estimator $\mathtt{EST}$ and $T\in\overline{\mathbb{N}}$, A continuous function $B:\mathcal{X}\times\Theta\to\mathbb{R}$ is an error-bound robust adaptive DT-CBF for (6) and $\mathtt{EST}$ if there exists $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$ such that, for all $x\in\mathcal{X}$, there exists $u\in\mathcal{U}$ satisfying
$\Delta B(x,u)-L_{B,x}\overline{w}-E^{\star}_{\eta}(x)\\ \geq-\alpha\left(B^{\star}(x)-\frac{\|\eta_{\theta,0}\|^{2}}{2\lambda_{\min}(\Gamma)}\right),\vskip-5.69046pt$ (34)
where $\Delta B(x,u)$ and $B^{\star}(x)$ are defined as in Definition 3.2, and $E^{\star}_{\eta}(x)\hskip-2.27626pt=\hskip-2.27626pt\left(L_{B,x}\|\bm{\phi}(x)\|+L_{B,\theta}\right)\|\eta_{\theta,0}\|$ with $L_{B,x}$ and $L_{B,\theta}$ given in Assumption 1. In this case, the condition (34) is called the error-bound robust adaptive CBC against disturbances and parametric model mismatch.
Following Theorem 2, a modified safety guarantee that uses Definition B.1 is given as follows:
Theorem 4.
Let Assumption 1 hold and assume $x_{0}\in\mathcal{S}_{\mathrm{ra},0}$. Given an estimator $\mathtt{EST}$ and $T\in\overline{\mathbb{N}}$, if $B$ is an error-bound robust adaptive DT-CBF for (6) and $\mathtt{EST}$ with its associated $\alpha\in\mathcal{K}_{\mathrm{e},\infty}\cap\mathcal{C}_{\mathrm{SL}}$, then for the error-bound robust adaptive online input condition
$B\big(f(x,u;\hat{\theta}_{t}),\hat{\theta}_{t}\big)-B(x,\hat{\theta}_{t})-L_{B,x}\overline{w}\\ -E_{\eta,t}(x)\geq-\alpha\left(B(x,\hat{\theta}_{t})-\frac{\|\eta_{\theta,t}\|^{2}}{2\lambda_{\min}(\Gamma)}\right),\vskip-5.69046pt$ (35)
the set $\overline{\mathcal{U}}_{\mathrm{ras},t}(x)$ defined as
[…]
is non-empty, $\forall x\in\mathcal{X}$ and $t\in\mathbb{I}_{[0,T-1]}$. The term $E_{\eta,t}(x)$ in (17) equals $L_{B,x}\|\bm{\phi}(x)\|\|\eta_{\theta,t}\|+L_{B,\theta}\|\delta_{\eta,t}\|$ with $\delta_{\eta,t}=\eta_{\theta,t+1}-\eta_{\theta,t}$. In addition, if time-varying state-feedback control policies $\overline{\pi}_{t}:\mathcal{X}\to\mathbb{R}^{m},t\in\mathbb{I}_{[0,T-1]}$ satisfies $\overline{\pi}_{t}(x)\in\overline{\mathcal{U}}_{\mathrm{ras},t}(x)$ for all $x\in\mathcal{X}$, then the time-varying closed-loop system
$x_{t+1}=f(x_{t},\overline{\pi}_{t}(x_{t});\theta^{\ast})+w_{t}\vskip-5.69046pt$ (37)
is robustly safe w.r.t. $\{\overline{\mathcal{S}}_{\mathrm{ra},t}\}^{T}_{t=0}$ and thus robustly safe w.r.t. $\{\mathcal{S}_{B,\hat{\theta}_{t}}\}^{T}_{t=0}$ (cf. Definition 3.1) against both parametric model mismatch and disturbances.
The proof of Theorem 4 is similar to that of Theorem 2 in Appendix A and is therefore omitted for brevity.
Remark B.2.
The advantage of using $\eta_{\theta,t}$ as in (32) is that $\overline{B}_{\mathrm{ra},t}$ can be explicitly determined. Moreover, using set-membership identification, it holds that $\eta_{\theta,t+1}\leq\eta_{\theta,t}$ (i.e., the error decays monotonically) since $\Theta_{t+1}\subseteq\Theta_{t}$, leading to $\overline{\mathcal{S}}_{\mathrm{ra},t}\subseteq\overline{\mathcal{S}}_{\mathrm{ra},t+1}$, i.e., the robustified safe set is monotonically expanding.
Appendix C More on Inherent Robustness
We first prove Theorem 3 using $V_{t}$ defined in (25). The proof is inspired by the proof of [32, Theorem B.13].
{pf}
If there exists a $t_{0}\in\mathbb{N}$ such that $x_{t_{0}}\in\mathcal{S}_{\mathrm{ra},t_{0}}$, then $x_{t}\in\mathcal{S}_{\mathrm{ra},t}$ holds for all $t\in\mathbb{I}_{[t_{0}:\infty]}$ by Theorem 2 since $\{B_{\mathrm{ra},t}\}^{\infty}_{t=0}$ is sequentially positive invariant for the system (6) with $u_{t}\in\mathcal{U}_{\mathrm{ras},t}(x_{t};p)$. Thus, in this case, $\lim_{t\to\infty}\mathrm{dist}_{\mathcal{S}_{\mathrm{ra},t}}(x_{t})=0$. On the other hand, assume $x_{t}\in\mathcal{X}\setminus\mathcal{S}_{\mathrm{ra},t}$; then we have
[…]
where the last inequality is due to $B_{\mathrm{ra},t}(x_{t})<0$ for $x_{t}\in\mathcal{X}\setminus\mathcal{S}_{\mathrm{ra},t}$ and $\alpha\in\mathcal{K}_{\mathrm{e},\infty}$. Consider the sequence $\{V_{t}(x_{t})\}^{\infty}_{t=0}$; given that $V_{t}(x_{t})\geq 0$ and (38), by the monotone convergence theorem [33], there exists a $V^{\ast}\geq 0$ such that $\lim_{t\to\infty}V_{t}(x_{t})=V^{\ast}$. Equivalently, $\forall\epsilon>0$, $\exists T^{\ast}(\epsilon)$ such that $\forall t\geq T^{\ast}(\epsilon)$, $V_{t}(x_{t})-V^{\ast}<\epsilon$. Next, we prove that $V^{\ast}=0$ by contradiction. Assume $V^{\ast}>0$, and let $\epsilon=-\alpha(-V^{\ast})$. Then, for any $t\geq T^{\ast}(-\alpha(-V^{\ast}))$, it holds that (i) $V^{\ast}\leq V_{t}(x_{t})<V^{\ast}-\alpha(-V^{\ast})$ and (ii) $V^{\ast}\leq V_{t+1}(x_{t+1})<V^{\ast}-\alpha(-V^{\ast})$, implying $V_{t+1}(x_{t+1})-V_{t}(x_{t})>\alpha(-V^{\ast})$. On the other hand, by (38), we have $V_{t+1}(x_{t+1})-V_{t}(x_{t})\leq\alpha(-V^{\ast})$, which leads to a contradiction. Therefore, $\lim_{t\to\infty}V_{t}(x_{t})=V^{\ast}=0$, implying $\lim_{t\to\infty}\mathrm{dist}_{\mathcal{S}_{\mathrm{ra},t}}(x_{t})=0$ by the continuity of $V_{t}$.   ∎
In fact, following the discussion in Appendix B, by constructing a sequence of energy functions using $\overline{\mathcal{S}}_{\mathrm{ra},t}$, it can been shown that the safe control input characterized by the certificate (34) also has the inherent robustness property, which is summarized in the following theorem:
Theorem 5.
Let Assumption 1 hold. If $B:\mathcal{X}\times\Theta\to\mathbb{R}$ is an error-bound robust adaptive DT-CBF for (6) and $\mathtt{EST}$ (cf. Definition B.1), then $x_{t}$ converges to $\{\overline{\mathcal{S}}_{\mathrm{ra},t}\}^{\infty}_{t=0}$ asymptotically (cf. Definition 5.1) on $\mathcal{X}$ for the system (6) with $u_{t}\in\overline{\mathcal{U}}_{\mathrm{ras},t}(x_{t})$.
Theorem 5 implies that $x_{t}$ will be regulated towards the safe sets $\{\overline{\mathcal{S}}_{\mathrm{ra},t}\}^{\infty}_{t=0}$ asymptotically on $\mathcal{X}$ for the closed-loop system (37). The proof of Theorem 5 closely follows that of Theorem 3 given above and is therefore omitted for brevity.
Biographies
Changrui Liu received the BSc degree (cum laude) from Harbin Institute of Technology (HIT) in Harbin, China, he graduated from Honors School, HIT and was awarded Best 10 Graduate, 2019. He then received the MSc degree from TU Delft, Aerospace Engineering. He is currently pursuing the PhD degree at the Delft Center for Systems and Control, TU Delft. His research interests are performance guarantees in machine learning and optimization.
Anil Alan received the B.Sc. degree from Mechanical Engineering Departments of Middle East Technical University, Ankara, Türkiye, in 2012, the M.Sc. degree from Bilkent University, Ankara, in 2017, and the Ph.D. degree from the University of Michigan, Ann Arbor, MI USA, in 2024. He is currently a Postdoc with the Delft University of Technology, Delft, The Netherlands. His research focuses on nonlinear safety-critical control with application to connected automated vehicles.
Shengling Shi is an assistant professor at Delft Center for Systems and Control, TU Delft, and he was a postdoctoral associate at the Massachusetts Institute of Technology, USA. He was also a postdoc at the Delft Center for Systems and Control, TU Delft, the Netherlands. He received the Ph.D. degree from the Eindhoven University of Technology in 2021. His research interests include system identification, model predictive control, and their applications.
Bart De Schutter (IEEE member since 2008, senior member since 2010, and fellow since 2019) is a full professor and head of department at the Delft Center for Systems and Control of Delft University of Technology in Delft, The Netherlands. He is also an IFAC fellow.